### v4

# Kalb-Ramond field, black holes and black strings in (2 + 1)D

Kalb-Ramond field, black holes and black strings in (2 + 1)D
Meseret Asrat
International Center for Theoretical Sciences
Tata Institute of Fundamental Research
Bengaluru, KA 560089, India
New rotating dilaton black hole and black string solutions in three spacetime dimensions are obtained. The background spacetime interpolates between Anti-de Sitter and a (an asymptotically) flat spacetime. The new black strings are characterized by their masses, angular momenta and axion charges. The uncharged black string solutions have only a single horizon. They contain a curvature singularity enclosed inside their horizons. Thus, they are also black hole solutions of Einstein-scalar gravity. On the other hand, the charged black string solutions have two horizons. They may or may not contain a curvature singularity depending on the ratio of their inner and outer horizons radii. When they contain a singularity, the singularity is either at or enclosed inside their inner horizons. We also obtain, in addition to the uncharged black strings, other new black hole solutions of Einstein-scalar gravity by applying duality transformation on the new charged black strings. The new black holes are described by their masses and angular momenta. We show that their angular momenta do not depend on their horizons radii. We also find that the masses of a class of the black holes can be made negative by adjusting, using the gauge ambiguity, the asymptotic value of the (independent component of the) Kalb-Ramond field in the dual black strings. We also discuss black hole and black string solutions with a curvature singularity at or beyond their outer or event horizons. Black hole solutions with a ring curvature singularity in between their inner and outer horizons are also presented.
09/24
1. Introduction
In this paper, we obtain new black hole and black string solutions in the string background obtained in [1]. The string background has the product topology ${\cal A}_{3}\times S^{3}\times{\cal X}_{4}$ where $S^{3}$ denotes a unit three-sphere and ${\cal X}_{4}$ is a compact four dimensional manifold. Examples include the Calabi-Yau manifolds $T^{4}$ and $K3$. The spacetime ${\cal A}_{3}$ interpolates between $AdS_{3}$ in the infrared and a (an asymptotically) linear dilaton spacetime ${\tenrm I\kern-1.79993ptR}\times S^{1}\times{\tenrm I\kern-1.79993ptR}$ in the ultraviolet. The worldsheet conformal field theory on ${\cal A}_{3}$ is an exact marginal current bilinear deformation of the worldsheet theory on $AdS_{3}$. In the boundary theory the deformation is equivalent to a deformation by an irrelevant operator of (left and right) dimensions $(2,2)$, see [[1]1,,  [2]2] and references therein. The resulting deformed theory is thus non-conformal and dual to string theory on ${\cal A}_{3}$ (times the internal product manifold ${\cal M}_{7}:=S^{3}\times{\cal X}_{4}$).
In the boundary theory the deformation coupling $\widehat{\gamma}$ has (left and right) mass dimensions $(-1,-1)$. The coupling can have either signs. In the case it is negative the deformation washes the holographic boundary away to infinity. On the other hand, in the case it is positive, in general it moves the boundary into the bulk. In Anti-de Sitter (AdS) spacetime moving the boundary into the bulk generates at onset a curvature singularity. For more and thorough discussions see [1]. Therefore, we will make in our following discussions a distinction between positive and negative coupling. Also since products of conformal field theories (CFTs) are still CFTs, we will simply ignore the (worldsheet) CFT on the internal manifold ${\cal M}_{7}$.
The string background ${\cal A}_{3}$ contains the Kalb-Ramond field. The field has only one independent component. At asymptotic infinity in general the component is a non-zero constant. We can freely set the constant to any arbitrary value using the gauge ambiguity. As we will show later in the paper, the usual gauge ambiguity in the definition of the Kalb-Ramond field or the constant value that its (independent) component takes at the boundary, i.e. at radial infinity, has a non-trivial consequence. In particular, it allows for black hole solutions (of Einstein-scalar gravity) with negative ADM (Arnowitt-Deser-Misner) masses.1 In a situation in which the metric (in the Einstein frame) is not asymptotically flat and/or the cosmological constant is not zero the usual ADM mass formula does not directly apply. In such a situation, the adapted or analogous ADM mass [3], however, is still given by the usual ADM expression except that now the reference metric is not flat and the lapse function is not unity. It equals the asymptotic value at radial infinity of the Brown-York (quasi-local) mass [4]. In this paper, we follow the Brown-York approach [5] and simply call the Brown-York mass at radial infinity the physical or ADM mass. See Appendix C for a review on the Brown-York energy and mass.
In three spacetime dimensions the Hodge dual of the Kalb-Ramond field is a (pseudo-)scalar which is identified as an axion field [6]. Its value at infinity gives an additional parameter to describe the solutions.2 T-duality relates it to angular momentum. The solutions we find are asymptotically flat (in the string frame) and they are exact solutions of classical string theory; we will be precise later in the paper.
The new black string solutions are parametrized by their masses, angular momenta and axion charges. The axion charge is the value of the (pseudo-)scalar at the boundary or spatial infinity. In general it is determined by the electric flux in the world-volume of the black string. We show that the axion charge is given only in terms of the deformation coupling. In particular it does not depend on the horizons (radii). The uncharged solutions, i.e., those black string solutions with zero axion charge, are also black hole solutions of Einstein-scalar gravity. We also obtain other new black hole solutions of Einstein-scalar gravity by applying duality transformation on the new charged black string solutions.3 See Appendix B for a review of the duality transformation. We show that their angular momenta are given by the dual black strings axion charges and thus do not depend on their horizons. We also show that the masses of a class of the black hole solutions can be made negative by changing (in the dual black strings) the value of the (component of the) Kalb-Ramond field at radial infinity.
The uncharged black strings have only a single horizon. They contain a curvature singularity with the structure of a ring. In the case the deformation coupling is negative the singularity is enclosed inside the event horizon.4 This particular class of solutions, i.e., the class of solutions characterized by zero axion charge and negative coupling (of $-1$), is already known in the literature [7]. It is obtained by directly solving the Einstein field equations with a scalar field. However, here we simply obtain it using duality and coordinate transformations from global $AdS_{3}$. We will discuss it in more detail later in the paper. Otherwise, i.e., for positive coupling, the singularity is beyond the event horizon. The charged black strings, on the other hand, possess two horizons. We show that depending on the ratio of the two horizons radii the solutions may or may not contain a curvature singularity. The singularity has the structure of a ring. In the case the deformation coupling is negative the singularity is at or enclosed inside the inner horizon.5 The inner (event) horizon is also referred to as Cauchy horizon. Otherwise, it is at or beyond the outer horizon. The new black hole solutions that we obtain by applying duality transformation have only a single horizon. They contain a curvature singularity. The singularity is enclosed inside the event horizon. To end, the black hole solutions that we will discuss, including those characterized by positive coupling, have only a single horizon and a ring singularity. The event horizon divides the spacetime into two regions. The interior region contains the singularity and the exterior region contains the rest of the spacetime. This will be our definition of interior and exterior. In the case the coupling is positive the ring singularity cannot be contracted to the (coordinate) origin without crossing the horizon(s).
The paper is organized as follows. In section two we review the rotating BTZ black hole [[8]8,,  [9]9]. In section three we obtain the new black hole and black string solutions. In section four we discuss the dilaton black hole solutions. These are black hole solutions of the Einstein-scalar gravity. We compute their ADM masses and angular momenta. In section five we summarize the main results and discuss future research directions. In Appendix A we review a parameterization of $AdS_{3}$ that in particular is better suited in relation to BTZ black hole. In Appendix B we review the target-space duality, also known as T-duality, transformation which we use in section three. In Appendix C we review the Brown-York quasi-local energy and mass. In Appendix D we review Kruskal approach which we use to remove coordinate singularities and obtain Penrose diagrams for the solutions we consider in the paper.
2. BTZ black hole
Unlike in higher dimensional gravity, in three dimensional gravity, the Riemann curvature tensor is determined completely by the Einstein tensor or equivalently the energy momentum tensor of the matter content.6 The Weyl curvature tensor is also identically zero. Thus, there are no physical excitations or propagating gravitons. Therefore, in the absence of matter the only solution is (locally) flat spacetime. However, in the case the cosmological constant is non-zero and negative, it has black hole solution [[8]8,,  [9]9]. The solution is referred to as BTZ black hole. It is a solution of Einstein equations in three spacetime dimensions with negative cosmological constant and without matter or fields.
The BTZ black hole is related to global $AdS_{3}$ (i.e., the universal cover of the $SU(1,1)\cong SL(2,{\tenrm I\kern-1.79993ptR})$ group manifold,) by a global coordinate transformation. $AdS_{3}$ (which here always denotes the infinitely-sheeted universal covering space of $SU(1,1)\cong SL(2,{\tenrm I\kern-1.79993ptR})$) is described in global coordinates by the metric
$ds^{2}=l^{2}(d\theta^{2}-{\tenrm cosh}^{2}\theta d\varphi^{2}+{\tenrm sinh}^{2}\theta d\psi^{2}),$ (2.1)2.1( 2.1 )
where $l$ is the radius of curvature and it is related to the (negative) cosmological constant $\Lambda$. $\psi$ is an angle variable and it has period $2\pi$. The time coordinate $\varphi$ takes its values in ${\tenrm I\kern-1.79993ptR}$ and thus it is not compact. The radial coordinate $\theta$ takes its values in $[0,\infty)$. We next discuss the coordinate transformation and obtain the BTZ black hole from $AdS_{3}$.
We first make the coordinate redefinition
$r=l{\tenrm sinh}\theta.$ (2.2)2.2( 2.2 )
The metric (2.1) becomes
$ds^{2}={dr^{2}\over{r^{2}\over l^{2}}+1}-\left(r^{2}+l^{2}\right)d\varphi^{2}+r^{2}d\psi^{2},$ (2.3)2.3( 2.3 )
We next make the coordinate transformation [[9]9[10]--[11]11]7 The complex number $i$ is related to a change of basis or coordinate patch, see Appendix A and [[12]12,,  [13]13].
$\matrix{\varphi=i\overline{\varphi},&\overline{\varphi}=-\left({\rho_{-}\over l^{2}}t-{\rho_{+}\over l}\omega\right),\cr\psi=i\overline{\psi},&\overline{\psi}=-\left({\rho_{+}\over l^{2}}t-{\rho_{-}\over l}\omega\right),}$ (2.4)2.4( 2.4 )
where $t,\omega\in{\tenrm I\kern-1.79993ptR}$, and extend the domain of $r$ to all the real numbers, i.e., $r={\overline{r}}\in{\tenrm I\kern-1.79993ptR}$, see A.5. We define, since we are in particular interested in the region exterior to the black hole horizons, $\rho$ as
${\overline{r}}^{2}=l^{2}\left({\rho^{2}-\rho_{+}^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}\right),$ (2.5)2.5( 2.5 )
where $\rho\geq\rho_{+}>\rho_{-}$. In our convention the coordinates $\rho$ and $t$ have mass dimension $-1$. Using the $GL(2)$ transformation (2.4) and the change of variable (2.5) in (2.3) we get the metric
$\eqalign{ds^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})\over l^{2}\rho^{2}}dt^{2}+\rho^{2}\left(d\omega-{\rho_{+}\rho_{-}\over l\rho^{2}}dt\right)^{2},\cr&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{(\rho^{2}-\rho_{+}^{2}-\rho^{2}_{-})\over l^{2}}dt^{2}-2{\rho_{+}\rho_{-}\over l}dtd\omega+\rho^{2}d\omega^{2}.}$ (2.6)2.6( 2.6 )
The metric is invariant under $\rho\to-\rho$. The BTZ black hole metric is obtained after identifying $\omega$ periodically. In the metric (2.6) we now identify $\omega$ with period $2\pi$, i.e., $\omega\sim\omega+2\pi$. This in turn implies $(\overline{\varphi},\overline{\psi})\sim(\overline{\varphi}+2\pi\rho_{+}/l,\overline{\psi}+2\pi\rho_{-}/l)$. Now $\rho$ is radial coordinate, i.e. $\rho\geq 0$. Therefore, BTZ black hole can be viewed as a quotient space of $AdS_{3}$. Thus, locally it is isometric to $AdS_{3}$. Furthermore, the scalar curvature is constant. It is given by8 In $AdS_{d}$ the scalar curvature and cosmological constant are given by $R=-{d(d-1)\over l^{2}},\quad\Lambda=-{(d-1)(d-2)\over 2l^{2}}.$ (2.7)2.7( 2.7 )
$R=-{6\over l^{2}}=6\cdot\Lambda,\quad\Lambda=-{1\over l^{2}},$ (2.8)2.8( 2.8 )
where $\Lambda$ is the cosmological constant. It approaches (without identifications) $AdS_{3}$ spacetime asymptotically.
The solution has two horizons. It has an event horizon at $\rho_{+}$ and an inner horizon at $\rho_{-}$. There is no curvature singularity inside the inner horizon. The spacetime region,
$\rho^{2}_{+}<\rho^{2}<\rho_{+}^{2}+\rho_{-}^{2},$ (2.9)2.9( 2.9 )
in which the Killing vector field $(\partial_{t})^{\mu}=(1,0,0):=\delta^{\mu}_{t}$ is space-like defines an ergosphere.9 The analytically extended BTZ metric is obtained by replacing $\rho$ with $i\rho$ or simply by setting $\rho^{2}=r\in{\tenrm I\kern-1.79993ptR}$. It contains closed time-like curves and thus usually it is not considered.
The BTZ black hole metric (2.6) satisfies the vacuum Einstein equations in $2+1$ spacetime dimensions,
$G_{ab}:=R_{ab}-{1\over 2}Rg_{ab}=-\Lambda g_{ab}={1\over l^{2}}g_{ab}.$ (2.10)2.10( 2.10 )
In an asymptotically flat spacetime the total mass (which equals the total energy since the lapse is one) is unambiguously defined. It is given by the ADM mass. The mass in $AdS$ and more generally in non-flat metrics is given by the usual ADM expression except that now the reference metric is not flat and the lapse is not one [[3]3,,[4]4]. It equals the Brown-York quasi-local mass [5] at radial infinity. The Brown-York mass is measured by a static observer. The angular momentum is given similarly by the Brown-York quasi-local angular momentum at radial infinity. It agrees with the Komar angular momentum [14], here adapted for $(2+1)$ dimensions.
The angular momentum and (analogous) ADM mass of the BTZ solution are given by10 In our convention, the mass $M$ is dimensionless. We set Newton’s gravitational constant $G_{N}$ to $1/8$ (in units of length). The angular momentum $J$ has mass dimension $-1$. For BTZ black hole the usual Komar mass formula adapted for (2 + 1) dimensions gives a divergent answer.
$J={2\rho_{+}\rho_{-}\over l},\quad M={\rho_{+}^{2}+\rho_{-}^{2}\over l^{2}}.$ (2.11)2.11( 2.11 )
In terms of $M$ and $J$ the BTZ (black hole) metric is
$ds^{2}={l^{2}\rho^{2}d\rho^{2}\over(\rho^{4}-l^{2}M\rho^{2}+{l^{2}J^{2}\over 4})}-{(\rho^{4}-l^{2}M\rho^{2}+{l^{2}J^{2}\over 4})\over l^{2}\rho^{2}}dt^{2}+\rho^{2}\left(d\omega-{J\over 2\rho^{2}}dt\right)^{2}.$ (2.12)2.12( 2.12 )
Thus, it is described by two parameters, its mass $M$ and angular momentum $J$. We will obtain $M$ and $J$ later in the paper following the Brown-York approach. In order for the solution to describe black hole or for a horizon to exist we assume
$M>0,\quad|J|\leq Ml.$ (2.13)2.13( 2.13 )
It is with these conditions that the BTZ metric contains black hole solutions.
We note that $(M,J)=(-1,0)$ is the $AdS_{3}$ metric (2.3). It is separated from the vacuum solution, i.e., $(M,J)=(0,0)$, or the continuous black hole spectrum, by a mass unit. The solutions with $-1<M<0,J=0$ have naked conical singularity at the origin $\rho=0$. For small $\rho$ the metric is
$ds^{2}=d{\widetilde{\rho}}^{2}-d\tau^{2}-M{\widetilde{\rho}}^{2}d\omega^{2},\quad\rho=\sqrt{-M}{\widetilde{\rho}},\quad\tau=\sqrt{-M}t.$ (2.14)2.14( 2.14 )
Thus, it describes point particle sources with negative cosmological constant [[15]15,,  [16]16].11 See [17] for numerical solutions of static non-singular vortices in $AdS_{3}$ that are not BTZ black holes. The zero mass black hole is identified in $(1,1)$ $AdS$ supergravity [18] as the Ramond sector ground state. The Neveu-Schwarz sector ground state is global $AdS_{3}$.
In the next section we obtain a family of rotating black hole and black string solutions by applying a similar $GL(2)$ transformation since in the infrared the spacetime ${\cal A}_{3}$ is well described by $AdS_{3}$. We explain this in the next section.
3. Rotating black hole and black string solutions
3.1. New black string solutions
We now consider the string background ${\cal A}_{3}$ obtained in [1]. ${\cal A}_{3}$ contains the metric $g_{\mu\nu}$, the Kalb-Ramond field $B_{\mu\nu}$ and the dilaton field $\Phi$. The string metric $g_{\mu\nu}$ is given by
$\eqalign{ds^{2}&=g_{ab}dx^{a}dx^{b},\cr&=l^{2}(d\theta^{2}-e^{2\phi}{\tenrm cosh}^{2}\theta d\varphi^{2}+e^{2\phi}{\tenrm sinh}^{2}\theta d\psi^{2}),}$ (3.1)3.1( 3.1 )
where
$e^{-2\phi}=1+\gamma^{2}-2\gamma{\tenrm cosh}(2\theta),$ (3.2)3.2( 3.2 )
and $\gamma$ is a dimensionless parameter. It is related to the deformation coupling $\widehat{\gamma}$ in the dual boundary field theory by $\widehat{\gamma}=\gamma R^{2}$ (up to a constant factor) here $R$ is the conformal radius of the boundary cylinder on which the undeformed CFT is defined.12 The conformal boundary of $AdS_{3}$ (i.e., the universal cover) is a time-like cylinder ${\tenrm I\kern-1.79993ptR}\times S^{1}$. As usual in AdS, we take $R=1$ in string units. The two form $B$ is given by
$B=B_{01}d\varphi\wedge d\psi,\quad B_{01}=-{1\over 2}l^{2}e^{2\phi}(\gamma-{\tenrm cosh}(2\theta)).$ (3.3)3.3( 3.3 )
The dilaton $\Phi$ is given by
$e^{2\Phi}=g_{s}^{2}|e^{2\phi}|={g_{s}^{2}\over\left|1+\gamma^{2}-2\gamma{\tenrm cosh}(2\theta)\right|},$ (3.4)3.4( 3.4 )
where $g_{s}$ is the string coupling at $\gamma=0$. Note the absolute value sign is necessary in the case $\gamma$ is positive to ensure the dilaton is real everywhere. The Kalb-Ramond field strength $H$ is given by
$\eqalign{H&=dB,\cr&=l^{2}(1-\gamma^{2}){\tenrm sinh}(2\theta)e^{4\phi}d\varphi\wedge d\psi\wedge d\theta.}$ (3.5)3.5( 3.5 )
The time coordinate $\varphi$ takes its value in ${\tenrm I\kern-1.79993ptR}$. The radial coordinates $\theta$ takes its value in ${\tenrm I\kern-1.79993ptR}^{+}$, i.e. $0\leq\theta<\infty$. The angular variable $\psi$ is periodic. It has period $2\pi(1-\gamma)$ to ensure that there is no conical singularity. The deformation coupling $\gamma$ takes the values in the range
$-1\leq\gamma\leq 1.$ (3.6)3.6( 3.6 )
Other values of $\gamma$ can be mapped onto the interval (3.6) by rescaling the coordinates and the string coupling at $\gamma=0$, i.e., $g_{s}$. The transformation rules under $\gamma\to 1/\gamma$ are $\varphi\to\varphi/\gamma$, $\psi\to-\psi/\gamma$ and $g_{s}^{2}\to g_{s}^{2}/\gamma^{2}$ (see Appendix E in [1]). The metric, the three form antisymmetric tensor and dilaton are invariant by these transformations.
$AdS_{3}$ (with Kalb-Ramond two form) is an exact solution to (classical) string theory, i.e., there are no $\alpha^{\prime}=l_{s}^{2}$ corrections to $AdS_{3}$. $l_{s}$ is the string length. The solution ${\cal A}_{3}$ is related to $AdS_{3}$ by T-dualities and coordinate shift (see [1] for details), and thus it is exact in $\alpha^{\prime}$. However, it is perturbative in $g_{s}$ [19]. On the world-sheet the deformation corresponds to an exact marginal deformation (see, e.g., [1]). More on this in a future work. Later in the section we also give previously known $\alpha^{\prime}$ exact (black string) solutions (in ${\cal A}_{3}$) as special cases.
In the low energy (super)gravity theory the metric, three form flux and dilaton are governed by the action
$S={1\over 2\kappa^{2}}\int d^{3}x\sqrt{-g}e^{-2\widetilde{\Phi}}\left[R+4\left(\nabla\widetilde{\Phi}\right)^{2}-{1\over 12}H^{2}-2\Lambda\right],$ (3.7)3.7( 3.7 )
where $x^{\mu}=(x^{0},x^{1},x^{2})=(\varphi,\psi,\theta)$ and
$e^{\widetilde{\Phi}}={e^{\Phi}\over g_{s}},\quad{1\over 2\kappa^{2}}={1\over 2\kappa_{0}^{2}g_{s}^{2}}={1\over 16\pi G_{N}},\quad\Lambda=-{2\over l^{2}},$ (3.8)3.8( 3.8 )
where $G_{N}$ is the gravitational coupling in three spacetime dimensions. The equations of motion that follow from the low energy (super)gravity action are
$\eqalign{R_{\mu\nu}+2\nabla_{\mu}\nabla_{\nu}\widetilde{\Phi}-{1\over 4}H_{\mu\rho\sigma}H_{\nu}^{\rho\sigma}=0,\cr\nabla^{\rho}\left(e^{-2\widetilde{\Phi}}H_{\mu\nu\rho}\right)=0,\cr R+4\nabla^{2}\widetilde{\Phi}-4\left(\nabla\widetilde{\Phi}\right)^{2}-{1\over 12}H^{2}-2\Lambda=0,}$ (3.9)3.9( 3.9 )
where $R=g^{\mu\nu}R_{\mu\nu}$ is the (string frame) Ricci scalar.
For $\gamma=0$, ${\cal A}_{3}$ is $AdS_{3}$. Assuming $\gamma$ is non-zero, for small $\theta$, i.e. ${\tenrm sinh}\theta\ll 1$, the metric (3.1) describes $AdS_{3}$ and for large $\theta$, i.e. ${\tenrm sinh}\theta\gg 1$ the metric is asymptotically flat. In what follows we simply apply the same $GL(2)$ transformation to obtain black hole and black string solutions in ${\cal A}_{3}$.
We now first make the coordinate redefinition
$r=l{\tenrm sinh}\theta.$ (3.10)3.10( 3.10 )
Using this the fields $g_{\mu\nu}$, $B_{\mu\nu}$ and $\Phi$ become
$\eqalign{ds^{2}&={dr^{2}\over{r^{2}\over l^{2}}+1}-e^{2\phi}\left(r^{2}+l^{2}\right)d\varphi^{2}+e^{2\phi}r^{2}d\psi^{2},\cr B&=B_{01}d\varphi\wedge d\psi,\quad B_{01}=-{1\over 2}l^{2}e^{2\phi}\left(\gamma-1-2{r^{2}\over l^{2}}\right),\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|,}$ (3.11)3.11( 3.11 )
where
$e^{-2\phi}=(1-\gamma)^{2}-4\gamma{r^{2}\over l^{2}}.$ (3.12)3.12( 3.12 )
We next make the coordinate transformation,
$\eqalign{\varphi&=-i\left({\rho_{-}\over l^{2}}t-{\rho_{+}\over l}x\right),\cr\psi&=-i\left({\rho_{+}\over l^{2}}t-{\rho_{-}\over l}x\right),}$ (3.13)3.13( 3.13 )
and the change of variable
$r^{2}=l^{2}\left({\rho^{2}-\rho_{+}^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}\right).$ (3.14)3.14( 3.14 )
The $GL(2)$ transformation (3.13) and the coordinate redefinition (3.14) give the black string solution described by the metric $g_{\mu\nu}$, Kalb-Ramond field $B_{\mu\nu}$ and dilaton field $\Phi$ which are given by
$\eqalign{ds^{2}&=g_{\mu\nu}dx^{\mu}dx^{\nu}\cr&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-e^{2\phi}{(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})\over l^{2}\rho^{2}}dt^{2}+e^{2\phi}\rho^{2}\left(dx-{\rho_{+}\rho_{-}\over l\rho^{2}}dt\right)^{2},\cr&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-e^{2\phi}{(\rho^{2}-\rho_{+}^{2}-\rho_{-}^{2})\over l^{2}}dt^{2}-2e^{2\phi}{\rho_{+}\rho_{-}\over l}dxdt+e^{2\phi}\rho^{2}dx^{2},\cr B&=B_{tx}dt\wedge dx,\quad B_{tx}={1\over 2}e^{2\phi}\left({\rho^{2}_{+}-\rho^{2}_{-}\over l}\right)\left[\gamma-1-2\left({\rho^{2}-\rho^{2}_{+}\over\rho^{2}_{+}-\rho^{2}_{-}}\right)\right],\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|.}$ (3.15)3.15( 3.15 )
where
$\eqalign{e^{-2\phi}&=(1-\gamma)^{2}-4\gamma\left({\rho^{2}-\rho^{2}_{+}\over\rho^{2}_{+}-\rho^{2}_{-}}\right),\cr&={\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}.}$ (3.16)3.16( 3.16 )
The time coordinate $t$ takes its value in ${\tenrm I\kern-1.79993ptR}$. We will in general assume $x$ is an angle coordinate. $\rho$ is a radial coordinate. We will in general assume, unless stated otherwise, that $\rho_{-}<\rho_{+}$. The periodicity of $x$ is chosen so that there is no conical singularity at the origin. We note that at $\rho=\infty$ the $B$ field, i.e., $B_{tx}$, is constant. We can adjust the constant using the gauge freedom. This has, as we will show in the next section, a non-trivial consequence. In particular, the constant enters into physical observables.
The three form flux is
$\eqalign{H&=dB,\cr&=H_{tx\rho}dt\wedge dx\wedge d\rho,}$ (3.17)3.17( 3.17 )
where
$\eqalign{H_{tx\rho}&={\partial B_{tx}\over\partial\rho},\cr&=-{2(1-\gamma^{2})\over l}\cdot{(\rho_{+}^{2}-\rho_{-}^{2})^{2}\rho\over\left[\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\right]^{2}},\cr&=-{2(1-\gamma^{2})\over l}\cdot e^{4\phi}\cdot\rho={2(1-\gamma^{2})\over l}\cdot\left({g\over\rho}\right),}$ (3.18)3.18( 3.18 )
where $g$ is the determinant of the string metric $g_{\mu\nu}$.
The black string solutions carry the axion charge (per unit length) $Q$ given by the value of the pseudo-scalar or axion field $e^{-2\Phi}{\star H}$ at the boundary or spatial infinity. It follows from the (super)gravity equations of motion, i.e., $\nabla^{\rho}(e^{-2\Phi}H_{\mu\nu\rho})=0$, that the axion is a constant and therefore $Q$ is simply given by
$Q=e^{-2\Phi}{\star H}=2\left({1-\gamma^{2}\over lg_{s}^{2}}\right),$ (3.19)3.19( 3.19 )
[…]
Note at $\gamma=0$ the metric describes BTZ black hole and it allows a non-zero Kalb-Ramond field strength $H$ [20]. The BTZ black hole solution (2.6) now with $\Phi=0$ and $B=B_{tx}dt\wedge dx$ where $B_{tx}=-\left[\rho^{2}-(\rho_{+}^{2}+\rho_{-}^{2})/2\right]/l$, i.e., the solution (3.15) at $\gamma=0$, is dual to the charged black string solution [20]14 In the appropriate gauge and with the appropriate coordinate changes., see also [[21]21,,  [22]22,,  [23]23,,  [24]24,,  [25]25] and [26]. We next consider different cases of the black string solution described by (3.15). We begin with the case where $\gamma$ is negative.15 The analytically extended solution, which is obtained by simply replacing $\rho$ by $i\widetilde{\rho}$ or $\rho^{2}$ by $r\in{\tenrm I\kern-1.79993ptR}$, in general contains closed time-like curves and thus we will not consider it. In the Penrose diagrams below we also do not include the regions which contain closed time-like curves.
3.2. Negative coupling
In this case we consider $-1<\gamma<0$. The scalar curvature is finite provided the ratio
${\rho_{-}\over\rho_{+}}<\mu:={1+\gamma\over 1-\gamma}.$ (3.20)3.20( 3.20 )
Thus, in the case $\rho_{-}/\rho_{+}<\mu$ we have rotating black string solutions with inner and outer horizons but no curvature singularity. The solutions are exact solutions of classical string theory since they are related to $AdS_{3}$ primarily by T-dualities (and coordinate transformations). $\rho_{-}$ denotes the inner horizon and $\rho_{+}$ denotes the outer horizon. The period of the angle $x$ is chosen to avoid a conical singularity at the origin. We identify
$x\sim x+2\pi\sqrt{{\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}\over\rho^{2}_{+}-\rho^{2}_{-}}}.$ (3.21)3.21( 3.21 )
The global structure of the solutions which satisfy the inequality (3.20) is described using Penrose diagram in Fig. 1. The solutions are axially symmetric. Therefore, we set the (shifted) angular coordinate (in each patch) to a particular value, see Appendix D for details. Thus, a point in the diagram represents a circle in spacetime, i.e. the orbit of the axial isometry. The diagram extends infinitely in both vertical directions.
Fig. 1: The plot depicts the global structure of the black string solutions with $\rho_{-}<\mu\cdot\rho_{+}$, where $\mu=(1+\gamma)/(1-\gamma),-1<\gamma<0$. The shifted angular coordinate is fixed in each of the patches to a particular value. $\rho_{\pm}$ represent horizons. The solutions have no singularity. Null geodesics are represented by straight lines at $\pi/4$ radians. In region II $\rho$ is time-like.
In the cases the condition (3.20) is not satisfied the scalar curvature is divergent at $\rho=\rho_{0}$ where
$\rho_{0}={1\over 2}\sqrt{{\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}\over\gamma}}.$ (3.22)3.22( 3.22 )
We note that at $\gamma=-1$ we have $\rho_{0}=\rho_{-}$. Thus, at $\gamma=-1$ the inner horizon becomes singular and the black string solution has only one event horizon. Since $\rho_{-}<\rho_{+}$, the curvature singularity is enclosed inside the event horizon. In general, for negative values of $\gamma$ and $\rho_{-}/\rho_{+}\geq\mu$, i.e., $-1\leq\gamma\leq(\rho_{-}-\rho_{+})/(\rho_{-}+\rho_{+})$, we have $0\leq\rho_{0}\leq\rho_{-}$. $\rho_{0}=0$ when $\gamma=(\rho_{-}-\rho_{+})/(\rho_{-}+\rho_{+})$. Thus, the singularity is at or enclosed inside the inner horizon.
We now consider the special cases $\rho_{0}=0$ and $\rho_{0}=\rho_{-}$. We first consider the case $\rho_{0}=0$. In this case we have
$\gamma={\rho_{-}-\rho_{+}\over\rho_{-}+\rho_{+}},\quad\rho_{+}>\rho_{-}.$ (3.23)3.23( 3.23 )
The black string solution (3.15) reduces to
$\eqalign{ds^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho^{2}_{-})}-{\rho_{+}^{2}\over(1-\gamma)^{2}}{(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho^{2}_{-})\over l^{2}\rho^{4}}dt^{2}+{\rho^{2}_{+}\over(1-\gamma)^{2}}\left(dx-{\rho_{+}\rho_{-}\over l\rho^{2}}dt\right)^{2},\cr B&=B_{tx}dt\wedge dx,\quad B_{tx}=-{1\over(1-\gamma)^{2}}{\rho_{+}^{2}\over l}\left(1-{\rho_{+}\rho_{-}\over\rho^{2}}\right),\quad Q=2\left({1-\gamma^{2}\over lg_{s}^{2}}\right),\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|={g_{s}^{2}\over(1-\gamma)^{2}}\left(\rho_{+}^{2}\over\rho^{2}\right).}$ (3.24)3.24( 3.24 )
The Ricci scalar is singular at the origin.16 Thus, here $x$ need not be periodic (with a definite period to avoid a conical singularity). Although we assumed $\rho_{+}>\rho_{-}$, i.e. $\gamma\neq 0$, to arrive at the solution, also here the limit $\rho_{+}\to\rho_{-}$ is mathematically well defined and gives non-trivial solution. In this limit, after an appropriate change of coordinates, we obtain $ds^{2}=(l^{2}/4)\left[dr^{2}/(r-r_{0})^{2}-d\tau^{2}(1-(r_{0}/r))^{2}+(d\chi-(r_{0}/r)d\tau)^{2}\right]$, $B=B_{\tau\chi}d\tau\wedge d\chi,B_{\tau\chi}=-(l^{2}/4)(1-(r_{0}/r))$, $e^{2\Phi}=g_{s}^{2}r_{0}/r$, where $r_{0}=\rho_{+}^{2}=\rho_{-}^{2}$. The scalar curvature is $R=\left[2(4r-7r_{0})r_{0}\right]/(l^{2}r^{2})$. Thus, $r=r_{0}$ is not a curvature singularity. We do not discuss this solution further here. We also note that both the dilaton and Kalb-Ramond fields are singular at the origin. We will discuss this family of solutions further later in the section. We in particular show that they are related to the solutions discussed in [21]. The global structure of the solutions for which $0\leq\rho_{0}<\rho_{-}$ is described using Penrose diagram in Fig. 2. The diagram is obtained in the same way as the example in Appendix D. In the diagram we set, in each patch, the (shifted) angular coordinate to a particular value since any other value is equivalent and related by axial symmetry. Therefore, a point in the diagram represents a circle in spacetime.
Fig. 2: The plot depicts the global structure of the black string solutions with $\rho_{-}\geq\mu\cdot\rho_{+}$, where $\mu=(1+\gamma)/(1-\gamma),-1<\gamma<0$. $\rho_{\pm}$ represent horizons. The solutions have a singularity at $\rho=\rho_{0}$. $\rho_{0}$ is represented by the zigzag lines. In general $0\leq\rho_{0}<\rho_{-}$. The singularity is time-like.
Now consider the case $\rho_{0}=\rho_{-}$, i.e., $\gamma=-1$. The black string solution (3.15) reduces at $\gamma=-1$ to
$\eqalign{ds^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{1\over 4}{(\rho^{2}-\rho_{+}^{2})(\rho_{+}^{2}-\rho_{-}^{2})\over l^{2}\rho^{2}}dt^{2}+{1\over 4}{\rho^{2}(\rho_{+}^{2}-\rho_{-}^{2})\over\rho^{2}-\rho_{-}^{2}}\left(dx-{\rho_{+}\rho_{-}\over l\rho^{2}}dt\right)^{2},\cr B_{tx}&=-{1\over 4}\left({\rho_{+}^{2}-\rho_{-}^{2}\over l}\right)={\tenrm const.},\quad Q=0,\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|,\quad e^{-2\phi}=4\left({\rho^{2}-\rho_{-}^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}\right).}$ (3.25)3.25( 3.25 )
The scalar curvature is given by
$R={4\over l^{2}}\left({\rho_{+}^{2}-\rho_{-}^{2}\over\rho^{2}-\rho_{-}^{2}}\right).$ (3.26)3.26( 3.26 )
Note the $B$ field, i.e. $B_{tx}$, is a non-zero constant. In general, using the gauge freedom we can set $B$ to zero. Thus, the solution describes a rotating dilaton black hole. This particular solution is already known in the literature [7]. It is obtained by directly solving the Einstein field equations with a scalar field. The novelty here, however, is that we simply obtained it using duality and coordinate transformations from global $AdS_{3}$ [1]. More on this black hole solution in the next section. The global structure of the solutions is described using Penrose diagram in Fig. 3. Each point represents a circle in spacetime.
Fig. 3: The plot depicts the global structure of the black hole solutions with $\gamma=-1$. $\rho_{+}$ represents the event horizon. The solutions have a space-like singularity at $\rho=\rho_{-}$. $\rho_{-}$ is represented by zigzag lines.
We note that in the case $\gamma=-1$ and $\rho_{-}=0$ the solution (3.15) (or either (3.24) or (3.25) with $\rho_{-}=0$) simplifies to
$\eqalign{ds^{2}&={l^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})}-{\rho_{+}^{2}(\rho^{2}-\rho_{+}^{2})\over 4l^{2}\rho^{2}}dt^{2}+{1\over 4}\rho^{2}_{+}dx^{2},\cr B&=B_{tx}dt\wedge dx,\quad B_{tx}=-{1\over 4}{\rho_{+}^{2}\over l}={\tenrm const.},\quad Q=0,\cr e^{2\Phi}&={g_{s}^{2}\over 4}\left({\rho^{2}_{+}\over\rho^{2}}\right).}$ (3.27)3.27( 3.27 )
The curvature singularity is at the origin. This solution in the gauge in which $B=0$ is dual to the non-rotating BTZ black hole with a non-vanishing $B$ [20], see also [22].17 As we will show later in the paper, for a general gauge, it is also equivalent to a non-rotating black hole in a spacetime with non-vanishing angular velocity at radial infinity and no $B$ field. It is the product of the two dimensional black hole [[27]27,,  [28]28] with $S^{1}$. To see this define
${\rho\over\rho_{+}}={\tenrm cosh}\theta\geq 1.$ (3.28)3.28( 3.28 )
The metric (3.27) becomes
$ds^{2}=l^{2}(d\theta^{2}-\tanh^{2}\theta d\tau^{2}+d\psi^{2}),$ (3.29)3.29( 3.29 )
where
$\tau:={\rho_{+}\over 2l}t,\quad\psi:={\rho_{+}\over 2}x.$ (3.30)3.30( 3.30 )
The dilaton is
$e^{-2\Phi}={4\over g_{s}^{2}}{\tenrm cosh}^{2}\theta.$ (3.31)3.31( 3.31 )
The scalar curvature is
$R={4\over l^{2}}{\tenrm sech}^{2}\theta.$ (3.32)3.32( 3.32 )
The analytic continuation $\tau\to i\tau$ and identification $\tau\sim\tau+2\pi$ give a metric which describes a manifold with the topology of a semi-cigar times a circle.
3.3. Positive coupling
In the case $0<\gamma\leq 1$ the scalar curvature is always divergent at $\rho=\rho_{0}$ (3.22), see (3.16). For $\gamma$ arbitrarily close to zero and positive, the metric describes BTZ black hole with a ring singularity at $\rho_{0}=\infty$. Therefore, from the onset the deformation develops curvature singularity at the boundary [1]. In general for positive $\gamma$, $\rho_{0}\geq\rho_{+}$. Thus, the singularity is either at or beyond the outer horizon. Therefore, it cannot be continuously contracted or deformed to the origin without crossing the horizon(s). Since we are assuming that $x$ is compact, the solution has closed time-like curves behind the singularity in the interior region, i.e., $\rho>\rho_{0}$. The Penrose diagram for $\rho_{0}>\rho_{+}$ is given in Fig. 4.
Fig. 4: The plot depicts the global structure of the black string solutions with $10<\gamma<$ 1. $\rho_{\pm}$ represent the horizons. The solutions have a time-like singularity at $\rho=\rho_{0}$. $\rho_{0}$ is represented by zigzag lines. For arbitrarily small positive $\gamma$, the singularity is at arbitrarily large $\rho=\rho_{0}$.
We note that at $\gamma=1$ we have $\rho_{0}=\rho_{+}$ and the black string solution (3.15) reduces to
$\eqalign{ds^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho_{+}^{2}-\rho^{2})(\rho_{-}^{2}-\rho^{2})}-{(\rho^{2}_{+}-\rho^{2}_{-})(\rho^{2}_{-}-\rho^{2})dt^{2}\over 4l^{2}\rho^{2}}+{(\rho_{+}^{2}-\rho_{-}^{2})\rho^{2}\over 4(\rho^{2}_{+}-\rho^{2})}\left(dx-{\rho_{-}\rho_{+}\over l\rho^{2}}dt\right)^{2},\cr B&=B_{tx}dt\wedge dx,\quad B_{tx}={\rho_{+}^{2}-\rho_{-}^{2}\over 4l}={\tenrm const.},\quad Q=0,\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|,\quad e^{-2\phi}=4\left({\rho_{+}^{2}-\rho^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}\right).}$ (3.33)3.33( 3.33 )
The scalar curvature is
$R={4\over l^{2}}\left({\rho_{+}^{2}-\rho_{-}^{2}\over\rho^{2}-\rho_{+}^{2}}\right).$ (3.34)3.34( 3.34 )
Thus, for an observer residing in the spacetime region $\rho<\rho_{-}$, i.e., the exterior region, the solution (3.33) has a singularity at $\rho=\rho_{0}=\rho_{+}$ beyond the horizon, i.e., $\rho_{0}>\rho_{-}$. Thus, in view of this observer, (3.33) describes a black hole. Since $x$ is compact, the solution contains closed time-like curves in the region $\rho>\rho_{0}=\rho_{+}$. Thus, they are located in the interior region. We note that the $B$ field, i.e. $B_{tx}$, is a non-zero constant. We can fix the $B$ field to zero using the gauge freedom. The Penrose diagram is given in Fig. 5.
Fig. 5: The plot depicts the global structure of the black string solution with $\gamma=1$. $\rho_{-}$ represents the horizon. The solution has a space-like singularity at $\rho_{0}=\rho_{+}$. $\rho_{0}$ is represented by zigzag lines.
In this paper, we will mainly focus on the black string solutions with negative $\gamma$. In what follows, we apply duality transformation on the black string solutions described by (3.15) to obtain other black hole solutions. See Appendix B for a review of the duality transformation.
3.4. New black hole solutions
We first boost the black string solutions (3.15) along the $x$ direction. To this end, we unwrap the coordinate $x$ and write
$\eqalign{t&=\widehat{t}{\tenrm cosh}(\alpha)+l\widehat{x}{\tenrm sinh}(\alpha),\cr lx&=l\widehat{x}{\tenrm cosh}(\alpha)+\widehat{t}{\tenrm sinh}(\alpha),}$ (3.35)3.35( 3.35 )
where $\alpha$ is given by
$\tanh(2\alpha)={2\rho_{+}\rho_{-}\over\rho_{+}^{2}+\rho_{-}^{2}}.$ (3.36)3.36( 3.36 )
The black string solutions now take the form
$\eqalign{d{\widehat{s}}^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-l^{-2}e^{2\phi}(\rho^{2}-\rho_{+}^{2})d{\widehat{t}}^{2}+e^{2\phi}(\rho^{2}-\rho_{-}^{2})d{\widehat{x}}^{2},\cr B&=B_{{\widehat{t}}{\widehat{x}}}d{\widehat{t}}\wedge d{\widehat{x}},\cr B_{{\widehat{t}}{\widehat{x}}}&={1\over 2}e^{2\phi}\left({\rho_{+}^{2}-\rho_{-}^{2}\over l}\right)\left[\gamma-1-2\left({\rho^{2}-\rho_{+}^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}\right)\right],\cr&={1\over 2l}e^{2\phi}\left[\rho_{+}^{2}(1+\gamma)+\rho_{-}^{2}(1-\gamma)-2\rho^{2}\right],\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|,}$ (3.37)3.37( 3.37 )
where
$e^{-2\phi}={\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}.$ (3.38)3.38( 3.38 )
We in particular note that there is no cross term in the metric.
Before we proceed further, we consider the special case in which
$\gamma={\rho_{-}-\rho_{+}\over\rho_{-}+\rho_{+}},\quad\rho_{+}>\rho_{-}.$ (3.39)3.39( 3.39 )
The black string solutions (3.37) reduce to
$\eqalign{d{\widehat{s}}^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{\rho_{+}^{2}\over l^{2}(1-\gamma)^{2}}\left(1-{\rho_{+}^{2}\over\rho^{2}}\right)d{\widehat{t}}^{2}+{\rho_{+}^{2}\over(1-\gamma)^{2}}\left(1-{\rho_{-}^{2}\over\rho^{2}}\right)d{\widehat{x}}^{2},\cr B&=B_{{\widehat{t}}{\widehat{x}}}d{\widehat{t}}\wedge d{\widehat{x}},\quad B_{{\widehat{t}}{\widehat{x}}}={\rho^{2}_{+}\over l(1-\gamma)^{2}}\left({\rho_{-}\rho_{+}\over\rho^{2}}-1\right),\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|={g_{s}^{2}\over(1-\gamma)^{2}}{\rho_{+}^{2}\over\rho^{2}}.}$ (3.40)3.40( 3.40 )
These particular non-rotating solutions are already known in the literature and they are first obtained in [21]. They are exact solutions of classical string theory. This family of solutions is related to (3.24) by the Lorentz boost (3.35). To put them into the form given in [21] we introduce the new coordinates
${\rho_{+}\over l(1-\gamma)}{\widehat{t}}={l\over 2}\tau,\quad{\rho_{+}\over(1-\gamma)}{\widehat{x}}={l\over 2}\chi,\quad\rho^{2}=r.$ (3.41)3.41( 3.41 )
This gives [21]
$\eqalign{d{\overline{s}}^{2}&={l^{2}\over 4}\left[{dr^{2}\over r^{2}\left(1-{r_{+}\over r}\right)\left(1-{r_{-}\over r}\right)}-\left(1-{r_{+}\over r}\right)d{\tau}^{2}+\left(1-{r_{-}\over r}\right)d{\chi}^{2}\right],\cr B&=B_{{\tau}{\chi}}d{\tau}\wedge d{\chi},\quad B_{{\tau}{\chi}}={l^{2}\over 4}\left({\sqrt{r_{-}r_{+}}\over r}-1\right),\cr-2\Phi&=\ln\left({|r|\over r_{+}}\right)+{\tenrm const.},}$ (3.42)3.42( 3.42 )
where $r_{\pm}=\rho_{\pm}^{2}$.18 The constant is ${\tenrm const.}=\ln\left[(1-\gamma)^{2}/g_{s}^{2}\right]$. It is invariant under $\gamma\to 1/\gamma$. Note $\Phi_{\tenrm there}=-2\Phi_{\tenrm here}$. Note also in general $\chi$ need not be periodic (with a definite period) and $r$ can take negative values. However here we will assume, for simplicity, $\chi$ is compact and $r$ is positive. The Ricci scalar is singular at $r=0$ and it is finite at $r=r_{\pm}$. We note that inside the event horizon, i.e., in the region $0<r<r_{+}$, the coordinate $\tau$ is everywhere space-like. We also note that $\chi$ becomes time-like in the region $0<r<r_{-}$. We will not discuss (3.42) further here. We refer to [21] for more details and a comprehensive discussion.
We first note that for $\gamma\neq 0$ the $B$ field, i.e. its component $B_{{\widehat{t}}{\widehat{x}}}$, is a constant at $\rho=\infty$. In general, we can use the gauge ambiguity in defining $B$ to set the constant at infinity to any arbitrary value. An equivalent Kalb-Ramond field, i.e., that gives the same $H=dB$, is
$\eqalign{{\widetilde{B}}_{{\widehat{t}}{\widehat{x}}}&=B_{{\widehat{t}}{\widehat{x}}}+{\lambda(\rho_{+}^{2}-\rho_{-}^{2})\over 2l},\cr&={1\over 2l}e^{2\phi}\left[\rho_{+}^{2}(1+\gamma)+\rho_{-}^{2}(1-\gamma)-2\rho^{2}\right]+{\lambda(\rho_{+}^{2}-\rho_{-}^{2})\over 2l},\cr&={1\over 2l}e^{2\phi}\left[\rho_{+}^{2}(1+\gamma)(1+\lambda(1+\gamma))+\rho_{-}^{2}(1-\gamma)(1-\lambda(1-\gamma))-2\rho^{2}(1+2\lambda\gamma)\right],\cr&={1\over 2l}e^{2\phi}\left[\rho_{+}^{2}(1+\lambda+\gamma(1+2\lambda)+\lambda\gamma^{2})+\rho_{-}^{2}(1-\lambda-\gamma(1-2\lambda)-\lambda\gamma^{2})-2\rho^{2}(1+2\lambda\gamma)\right].}$ (3.43)3.43( 3.43 )
Under $\gamma\to 1/\gamma$, $\lambda$ transforms as $\lambda\to-\gamma(1+\gamma\lambda)$ so that not only $H$ but ${\widetilde{B}}$ is also invariant.
We next periodically identify ${\widehat{x}}$ and perform $T$-duality. This gives the dilaton black hole solution described by the metric and dilaton,
$\eqalign{d{\widetilde{s}}^{2}&=g_{\rho\rho}d\rho^{2}+{1\over g_{{\widehat{x}}{\widehat{x}}}}(d{\widetilde{x}}-{\widetilde{B}}_{{\widehat{t}}{\widehat{x}}}d{\widehat{t}})^{2}+g_{{\widehat{t}}{\widehat{t}}}d{\widehat{t}}^{2},\cr&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{+}^{2})\over l^{2}\left[\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\right]}d{\widehat{t}}^{2}\cr&+\left[{\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\over(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{-}^{2})}\right]\cr&\cdot\left\{d{\widetilde{x}}-{(\rho_{+}^{2}-\rho_{-}^{2})\left[\rho_{+}^{2}(1+\gamma)(1+\lambda(1+\gamma))+\rho_{-}^{2}(1-\gamma)(1-\lambda(1-\gamma))-2\rho^{2}(1+2\lambda\gamma)\right]\over 2l\left[\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\right]}d{\widehat{t}}\right\}^{2},\cr&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{(\rho_{+}^{2}-\rho_{-}^{2})(4\lambda(1+\gamma\lambda)\rho^{2}+\rho_{-}^{2}(1-(1-\gamma)\lambda)^{2}-\rho_{+}^{2}(1+(1+\gamma)\lambda)^{2})\over 4l^{2}(\rho^{2}-\rho_{-}^{2})}d{\widehat{t}}^{2}\cr&-2{\rho_{-}^{2}(1-\gamma)(1-(1-\gamma)\lambda)+\rho_{+}^{2}(1+\gamma)(1+(1+\gamma)\lambda)-2(1+2\gamma\lambda)\rho^{2}\over 2l(\rho^{2}-\rho_{-}^{2})}d{\widetilde{x}}d{\widehat{t}}\cr&+{\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\over(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{-}^{2})}d{\widetilde{x}}^{2},\cr e^{-2\widetilde{\Phi}}&=|g_{s}^{-2}(\rho^{2}-\rho_{-}^{2})|,}$ (3.44)3.44( 3.44 )
where ${\widetilde{x}}$ is the dual coordinate. Under $\gamma\to 1/\gamma$, ${\widetilde{x}}\to-{\widetilde{x}}/\gamma$ and ${\widehat{t}}\to{\widehat{t}}/\gamma$. We note that the $\widetilde{B}$ field, and thus in turn the gauge parameter $\lambda$, enters into the coefficient of the cross term $d{\widehat{t}}d{\widetilde{x}}$. The solution is invariant under $\gamma\to 1/\gamma$ because $\lambda$ transforms as $\lambda\to-\gamma(1+\gamma\lambda)$. We will assume $0<\lambda<-1/\gamma$ so that the solution describes a black hole. The scalar curvature is given by
$R={4(\rho_{+}^{2}-\rho_{-}^{2})\over l^{2}(\rho^{2}-\rho_{-}^{2})}.$ (3.45)3.45( 3.45 )
Thus, the black hole solutions (labelled by the triplet $(\rho_{-},\rho_{+};\lambda)$) have a curvature singularity enclosed inside their event horizons.
We specially note that $R$ does not depend on the coupling $\gamma$. In particular for $\lambda=-1/(2\gamma)$ and $\gamma=(\rho_{-}-\rho_{+})/(\rho_{-}+\rho_{+})$ the solution can be put using the coordinates (3.41) into the form
$\eqalign{d{\widetilde{s}}^{2}&={l^{2}\over 4}\left[{dr^{2}\over\left(1-{r_{+}\over r}\right)\left(1-{r_{-}\over r}\right)r^{2}}-\left(1-{r_{+}\over r}\right)d\tau^{2}+{1\over\left(1-{r_{-}\over r}\right)}\left(d\chi-{\sqrt{r_{-}r_{+}}\over r}d\tau\right)^{2}\right],\cr-2\widetilde{\Phi}&=\ln(|r-r_{-}|)+{\tenrm const.},}$ (3.46)3.46( 3.46 )
where $r_{\pm}=\rho_{\pm}^{2}$. In the limit $r_{+}\to r_{-}:=r_{0}$ this simplifies to
$\eqalign{d{\widetilde{s}}^{2}&={l^{2}\over 4}\left[{dr^{2}\over\left(1-{r_{0}\over r}\right)^{2}r^{2}}-\left(1-{r_{0}\over r}\right)d\tau^{2}+{1\over\left(1-{r_{0}\over r}\right)}\left(d\chi-{r_{0}\over r}d\tau\right)^{2}\right],\cr-2\widetilde{\Phi}&=\ln(|r-r_{0}|)+{\tenrm const.}.}$ (3.47)3.47( 3.47 )
We note from (3.45) that in this limit the string frame Ricci scalar $R=0$. However, we also note that the dilaton diverges at $r=r_{0}$. In fact, in the Einstein frame, the scalar curvature depends on $r$ and it is singular at $r=r_{0}$. It is given by $R=-8/(l^{2}(r-r_{0})^{2})$. Thus, (in the Einstein frame) the solution (3.47) describes a naked singularity. In this paper we in general assume $\rho_{+}>\rho_{-}$.
We note that in the case $\gamma=-1$ the black hole solution (3.44) reduces to
$\eqalign{d{\widetilde{s}}^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{+}^{2})d{\widehat{t}}^{2}\over 4l^{2}(\rho^{2}-\rho_{-}^{2})}+{4\over(\rho_{+}^{2}-\rho_{-}^{2})}\left[d{\widetilde{x}}+{(\rho_{+}^{2}-\rho_{-}^{2})(1-2\lambda)\over 4l}d{\widehat{t}}\right]^{2},\cr e^{-2\widetilde{\Phi}}&=|g_{s}^{-2}(\rho^{2}-\rho_{-}^{2})|.}$ (3.48)3.48( 3.48 )
We note that for $\rho_{-}=0$ it is equivalent to the solution (3.27) in the sense both have a horizon at $\rho_{+}$ and a singularity at $\rho=\rho_{-}=0$. However, here the black hole is in a spacetime with non-vanishing angular velocity at radial infinity and no $B$ field.
The black hole solutions given in (3.44) are described by the four parameters $\rho_{-}$, $\rho_{+}$, $\gamma$ and $\lambda$. Thus, in general they are not identical. To the best of my knowledge the black hole solutions (3.44) have not been reported elsewhere in the literature. We next find the (analogous) ADM masses and angular momenta of the solutions (3.25) and (3.44). The ADM masses are given in a similar way as in the BTZ black hole by the asymptotic values of the Brown-York quasi-local masses at radial infinity. The angular momenta are given by the Brown-York quasi-local angular momenta which (as we momentarily show) are independent of $\rho$ and thus constant.
4. Dilatonic rotating black hole solutions
In this section we consider the black hole solutions (3.25) and (3.44). They are exact solutions of classical string theory, i.e., they have exact worldsheet CFT descriptions. The reason is essentially because we found them by applying T-dualities and coordinate transformations on $AdS_{3}$. More on the deformed CFTs in a future work. The solutions are again given below for convenience. The rotating black hole solution (3.25) is given by
$\eqalign{ds^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{1\over 4}{(\rho^{2}-\rho_{+}^{2})(\rho_{+}^{2}-\rho_{-}^{2})\over l^{2}\rho^{2}}dt^{2}+{1\over 4}{\rho^{2}(\rho_{+}^{2}-\rho_{-}^{2})\over\rho^{2}-\rho_{-}^{2}}\left(dx-{\rho_{+}\rho_{-}\over l\rho^{2}}dt\right)^{2},\cr&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{1\over 4l^{2}}{(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{+}^{2}-\rho_{-}^{2})\over\rho^{2}-\rho_{-}^{2}}dt^{2}-{1\over 2l}{\rho_{+}\rho_{-}(\rho_{+}^{2}-\rho_{-}^{2})\over\rho^{2}-\rho_{-}^{2}}dxdt\cr&+{1\over 4}{\rho^{2}(\rho_{+}^{2}-\rho_{-}^{2})\over\rho^{2}-\rho_{-}^{2}}dx^{2},\cr e^{-2\Phi}&=g_{s}^{-2}\left|e^{-2\phi}\right|,\quad e^{-2\phi}=4\left({\rho^{2}-\rho_{-}^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}\right).}$ (4.1)4.1( 4.1 )
The rotating black hole solution (3.44) is given by
$\eqalign{d{\widetilde{s}}^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{+}^{2})\over l^{2}\left[\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\right]}d{\widehat{t}}^{2}\cr&+\left[{\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\over(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{-}^{2})}\right]\cr&\cdot\left\{d{\widetilde{x}}-{(\rho_{+}^{2}-\rho_{-}^{2})\left[\rho_{+}^{2}(1+\gamma)(1+\lambda(1+\gamma))+\rho_{-}^{2}(1-\gamma)(1-\lambda(1-\gamma))-2\rho^{2}(1+2\lambda\gamma)\right]\over 2l\left[\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\right]}d{\widehat{t}}\right\}^{2},\cr e^{-2\widetilde{\Phi}}&=g_{s}^{-2}\left|e^{-2\phi}\right|,\quad e^{-2\phi}=(\rho^{2}-\rho_{-}^{2}).}$ (4.2)4.2( 4.2 )
The parameter $\lambda$ is constrained to lie within $0<\lambda<-{1/\gamma}$ so the solution describes a black hole.
The scalar curvature in both cases is given by
$R={4\over l^{2}}\left({\rho_{+}^{2}-\rho_{-}^{2}\over\rho^{2}-\rho_{-}^{2}}\right).$ (4.3)4.3( 4.3 )
Thus, the inner horizon disappears and becomes singular. This is related to the inner horizon in BTZ being unstable [29]. However, it is important to note, as we saw earlier, that the outer horizon can also instead be singular and thus unstable upon coupling the theory to matter. Note also, however, the singularity now is in the interior region, see the case $\gamma=1$ (3.33). We hope to study this phenomenon or scenario further in the future.
The solution (4.2) has similar Penrose diagram to that of (4.1). It is depicted in Fig. 3.
The solutions solve the (low energy) (super)gravity action [30]. In string frame it is simply given by setting $H=0$ in (3.7),
$S={1\over 2\kappa^{2}}\int d^{3}x\sqrt{-g}e^{-2\phi}\left[R+4(\nabla\phi)^{2}-2\Lambda\right].$ (4.4)4.4( 4.4 )
The equations of motion are
$\eqalign{R_{\mu\nu}+2\nabla_{\mu}\nabla_{\nu}\phi&=0,\cr R+4\nabla^{2}\phi-4(\nabla\phi)^{2}-2\Lambda&=0.}$ (4.5)4.5( 4.5 )
In the Einstein frame the low energy action is
$S={1\over 2\kappa^{2}}\int d^{3}x\sqrt{-\widetilde{g}}\left[\widetilde{R}-4(\widetilde{\nabla}\phi)^{2}-V(\phi)\right],\quad V(\phi)=2\Lambda e^{4\phi},$ (4.6)4.6( 4.6 )
where $\widetilde{g}_{ab}$ is the metric in the Einstein frame and it is given by
$\widetilde{g}_{ab}=e^{-4\phi}g_{ab}.$ (4.7)4.7( 4.7 )
Note under the conformal rescaling $\widetilde{g}_{ab}=e^{2\alpha}g_{ab}$ the scalar curvature is not invariant. It transforms to
$\widetilde{R}=e^{-2\alpha}\left[R-2(d-1)\nabla^{2}\alpha-(d-2)(d-1)(\nabla\alpha)^{2}\right].$ (4.8)4.8( 4.8 )
Thus, for $\alpha=-2\phi$ and $d=3$ we find
$\eqalign{\widetilde{R}&=e^{4\phi}\left[R-4\nabla^{2}\alpha-2(\nabla\alpha)^{2}\right],\cr&=-{(\rho_{+}^{2}-\rho_{-}^{2})^{2}\over 4l^{2}}{\left[\rho_{+}^{2}-\rho_{-}^{2}+2(\rho^{2}-\rho_{-}^{2})\right]\over(\rho^{2}-\rho_{-}^{2})^{3}}.}$ (4.9)4.9( 4.9 )
The equations of motion that we get from varying the action (4.6) are
$\eqalign{\widetilde{R}_{ab}&=4\widetilde{\nabla}_{a}\phi\widetilde{\nabla}_{b}\phi+\widetilde{g}_{ab}V(\phi),\cr\widetilde{\nabla}^{2}\phi&={1\over 8}{\partial V(\phi)\over\partial\phi}.}$ (4.10)4.10( 4.10 )
The first equation is the Einstein equations for the metric
$\widetilde{G}_{ab}:=\widetilde{R}_{ab}-{1\over 2}\widetilde{g}_{ab}\widetilde{R}=\kappa^{2}\widetilde{T}_{ab},$ (4.11)4.11( 4.11 )
where the energy-momentum tensor is
$\kappa^{2}\widetilde{T}_{ab}=4\widetilde{\nabla}_{a}\phi\widetilde{\nabla}_{b}\phi-2\widetilde{g}_{ab}(\widetilde{\nabla}\phi)^{2}-{1\over 2}\widetilde{g}_{ab}V(\phi).$ (4.12)4.12( 4.12 )
To compute the (analogous) ADM masses and (Komar) angular momenta we need the metrics in the Einstein frame.
In the Einstein frame, the string (frame) solution (4.1), is given by
$\eqalign{ds^{2}&={16l^{2}\over(\rho_{+}^{2}-\rho_{-}^{2})^{2}}{\rho^{2}(\rho^{2}-\rho_{-}^{2})\over(\rho^{2}-\rho_{+}^{2})}d\rho^{2}-{4\over l^{2}(\rho_{+}^{2}-\rho_{-}^{2})}(\rho^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{+}^{2}-\rho_{-}^{2})dt^{2}-{8\rho_{+}\rho_{-}(\rho^{2}-\rho_{-}^{2})\over l(\rho_{+}^{2}-\rho_{-}^{2})}dtdx\cr&+{4\rho^{2}(\rho^{2}-\rho_{-}^{2})\over(\rho_{+}^{2}-\rho_{-}^{2})}dx^{2},\cr&={16l^{2}\over(\rho_{+}^{2}-\rho_{-}^{2})^{2}}{\rho^{2}(\rho^{2}-\rho_{-}^{2})\over(\rho^{2}-\rho_{+}^{2})}d\rho^{2}-{4\over l^{2}(\rho_{+}^{2}-\rho_{-}^{2})}{(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})^{2}\over\rho^{2}}dt^{2}\cr&+{4\rho^{2}(\rho^{2}-\rho_{-}^{2})\over(\rho_{+}^{2}-\rho_{-}^{2})}\left(dx-{\rho_{+}\rho_{-}\over l\rho^{2}}dt\right)^{2},\cr e^{-2\Phi}&=\left|4g_{s}^{-2}\left({\rho^{2}-\rho_{-}^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}\right)\right|.}$ (4.13)4.13( 4.13 )
We can write the black hole solution (4.13) into the form
$\eqalign{d\widetilde{s}^{2}&=N^{2}(r)dr^{2}-K^{2}(r)dt^{2}+r^{2}\left(d\theta+Zdt\right)^{2},\cr e^{-2\widetilde{\Phi}}&=|g_{s}^{-2}f(r)|,}$ (4.14)4.14( 4.14 )
where $r$ is given by the equation
$\rho^{2}={1\over 2}\left(\rho_{-}^{2}+\sqrt{\rho_{-}^{4}+r^{2}(\rho_{+}^{2}-\rho_{-}^{2})}\right),$ (4.15)4.15( 4.15 )
and $\theta=x$. The singularity is at $r=0$ and the horizon is at $r=2\rho_{+}$. The metric components and dilaton $N,K,Z$ and $f$ are some known functions of $r$.
Similarly, the string frame solution (4.2), is given in the Einstein frame by
$\eqalign{d{\widetilde{s}}^{2}&=-{(\rho^{2}-\rho_{-}^{2})^{2}(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{+}^{2})\over l^{2}\left[\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\right]}dt^{2}+{l^{2}(\rho^{2}-\rho_{-}^{2})\rho^{2}\over(\rho^{2}-\rho_{+}^{2})}d\rho^{2}\cr&+\left({\rho^{2}-\rho_{-}^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}\right)\left[\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\right]\cr&\cdot\left\{dx-{(\rho_{+}^{2}-\rho_{-}^{2})\left[\rho_{+}^{2}(1+\gamma)(1+\lambda(1+\gamma))+\rho_{-}^{2}(1-\gamma)(1-\lambda(1-\gamma))-2\rho^{2}(1+2\lambda\gamma)\right]\over 2l\left[\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\right]}dt\right\}^{2},\cr e^{-2\widetilde{\Phi}}&=|g_{s}^{-2}(\rho^{2}-\rho_{-}^{2})|.}$ (4.16)4.16( 4.16 )
For later convenience, we set here ${\widetilde{x}}=x$ and ${\widehat{t}}=t$. We can also write the black hole solution (4.16) into the form (4.14). Here $r$ is given by the equation
$\rho^{2}=\rho_{-}^{2}-{1\over 8\gamma}\left[\sqrt{(\rho_{+}^{2}-\rho_{-}^{2})^{2}(1+\gamma)^{4}-16\gamma(\rho_{+}^{2}-\rho_{-}^{2})r^{2}}-(\rho_{+}^{2}-\rho_{-}^{2})(1+\gamma)^{2}\right],$ (4.17)4.17( 4.17 )
and $\theta=x$. The singularity is at $r=0$ and the horizon is at $r=(1-\gamma)\sqrt{\rho_{+}^{2}-\rho_{-}^{2}}$. The metric components and dilaton $N,K,Z$ and $f$ are some other known functions of $r$.
The black hole solution (4.13), as we mentioned earlier, is already known in the literature and was first obtained earlier in [7]. The authors [7] studied the Einstein-scalar gravity in three spacetime dimensions with the scalar in an exponential potential. The solution (4.13) belongs to the family of black hole solutions obtained in [7]. In [7] it corresponds to the case $N=1$. To write (4.13) in the form given in [7] we need to make the change of coordinate
$r^{2\nu}=4\left({\rho^{2}-\rho_{-}^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}\right),\quad\nu={1\over 4}.$ (4.18)4.18( 4.18 )
We now use the Einstein frame metrics to obtain the ADM masses and angular momenta of the rotating black hole solutions (4.13) and (4.16). We follow the Brown-York quasi-local approach [5]. We have reviewed the approach and summarized relevant formulae in Appendix C. Note that the dilaton is minimally coupled to gravity, see the action (4.6). It also only depends on $\rho$. Thus, it does not contribute to the ADM mass and angular momentum. See eq. (C.15) in the appendix and the discussion in the paragraph following it.
For a $(2+1)$ stationary and axisymmetric spacetime of the form [31]
$ds^{2}=-N^{2}dt^{2}+{1\over L^{2}}dr^{2}+K^{2}(Zdt+d\phi)^{2},$ (4.19)4.19( 4.19 )
the quasi-local energy $E$, angular momentum $j$ and mass $m$ are given by (with $\kappa^{2}=\pi$ in some length unit)
$\eqalign{j&={LZ^{\prime}K^{3}\over N},\cr E&=2(\varepsilon-LK^{\prime}),\cr m&=NE-jZ,}$ (4.20)4.20( 4.20 )
where $\varepsilon$ determines the zero of the energy. It is given by $\varepsilon=LK^{\prime}$ for a particular solution. Primes denote derivatives with respect to $r$. $j$ agrees with the usual Komar angular momentum formula (adapted for $(2+1)$ dimensions). For the derivation of (4.20) and review see the appendix. See also, e.g., [[7]7,,  [31]31,,  [32]32,,  [33]33].
The angular momentum $J$ and (analogous) ADM mass $M$ are then given by
$\eqalign{J&=\lim_{r\to\infty}j(r),\cr M&=\lim_{r\to\infty}m(r).}$ (4.21)4.21( 4.21 )
It is important to keep in mind that the mass $m$ is measured by the static observer whose four-velocity is $u^{\mu}_{(1)}=\delta^{\mu}_{t}=(1,0,0)$, i.e., tangent to the static Killing field $\xi^{\mu}=(\partial_{t})^{\mu}$. The energy $E$ is measured by the stationary observer whose four-velocity is $u^{\mu}_{(2)}=(1/N)\delta^{\mu}_{t}-(Z/N)\delta^{\mu}_{\phi}$.19 Note in general $u^{\mu}_{(2)}$ gives, in the large $r$ limit, the Kodama field [[34]34,,[35]35]. The Kodama vector field is given by $\eqalign{k&=k^{t}\partial_{t}+k^{\phi}\partial_{\phi},\cr k^{t}&={1\over N}LK^{\prime},\cr k^{\phi}&=-{Z\over N}LK^{\prime}+{1\over 2K^{2}}\cdot{LK^{3}Z^{\prime}\over N},\cr&=-{Z\over N}LK^{\prime}+{j\over 2K^{2}}.}$ (4.22)4.22( 4.22 ) In general, the Kodama field is not a Killing field. At large $r$ it is proportional to $Nu^{\mu}_{(2)}=(\partial_{t})^{\mu}-Z(\partial_{\phi})^{\mu}$. Its world-line is perpendicular to the $t={\tenrm constant}$ space-like hyper-surface. Thus, the energy and mass are equal only when the shift $Z=0$ and the lapse $N=1$.
In what follows we use (4.21) to obtain the angular momentum and ADM mass of the black hole solutions. We begin with the BTZ black hole. This helps to clarify the main points and illustrate the approach.
For the rotating BTZ black hole (2.6) we have
$\eqalign{N&=\sqrt{{(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})\over l^{2}\rho^{2}}},\cr L&=N,\cr K&=\rho,\cr Z&=-{\rho_{+}\rho_{-}\over l\rho^{2}},\cr\Lambda&=-{1\over l^{2}}.}$ (4.23)4.23( 4.23 )
We obtain the quasi-local angular momentum $j$
$j={2\rho_{+}\rho_{-}\over l}.$ (4.24)4.24( 4.24 )
Thus, the (Komar) angular momentum $J=j$. The quasi-local energy is
$E=-2\sqrt{{\rho^{2}\over l^{2}}+{J^{2}\over 4\rho^{2}}-{\rho^{2}_{+}+\rho_{-}^{2}\over l^{2}}}+2\varepsilon,\quad\rho\geq\rho_{+}.$ (4.25)4.25( 4.25 )
A particular solution to Einstein field equations is global $AdS_{3}$ (2.3) or the conformally flat spacetime
$\eqalign{ds^{2}&={l^{2}d\rho^{2}\over\rho^{2}}-{\rho^{2}\over l^{2}}d\varphi^{2}+\rho^{2}d\psi^{2},\cr\Lambda&=-{1\over l^{2}}.}$ (4.26)4.26( 4.26 )
This solution (4.26) gives $\varepsilon=\sqrt{-\Lambda\rho^{2}}$. The particular solution is chosen such that at constant time or $\varphi$ it asymptotically coincides or merges at large $\rho$ smoothly with the solution (2.6) [[3]3,,[4]4]. Thus, the quasi-local mass $m$ is given by20 Note here the boundary $B_{t}$ is at large $\rho$, see Appendix C. Thus, in the expressions for $m$ and $E$, $\rho$ is assumed to be large. The same applies for the quasi-local masses and energies that appear below. Here they serve as intermediate results.
$m=-2\left({\rho^{2}\over l^{2}}+{J^{2}\over 4\rho^{2}}-{\rho^{2}_{+}+\rho_{-}^{2}\over l^{2}}\right)+2\sqrt{{\rho^{2}\over l^{2}}}\sqrt{{\rho^{2}\over l^{2}}+{J^{2}\over 4\rho^{2}}-{\rho^{2}_{+}+\rho_{-}^{2}\over l^{2}}}+{J^{2}\over 2\rho^{2}}.$ (4.27)4.27( 4.27 )
The (analogous) ADM mass is then
$M=m(\rho\to\infty)={\rho^{2}_{+}+\rho_{-}^{2}\over l^{2}}.$ (4.28)4.28( 4.28 )
Note the quais-local mass $m$ depends on $\varepsilon$, i.e., the reference background. Different $\varepsilon$ gives different $m$. The global $AdS_{3}$ (2.3), for instance, gives $\varepsilon=\sqrt{-\Lambda\rho^{2}+1}$. However, the ADM mass $M$ is independent of the reference solution or regularization and it is finite.
We now consider the black hole metric (4.13). The (analogous) ADM mass and (Komar) angular momentum are obtained in [7], however, for convenience and completeness we obtain them here again. We have
$\eqalign{N&=\sqrt{{4\over l^{2}(\rho_{+}^{2}-\rho_{-}^{2})}{(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})^{2}\over\rho^{2}}},\cr L&=\sqrt{{(\rho_{+}^{2}-\rho_{-}^{2})^{2}\over 16l^{2}}{(\rho^{2}-\rho_{+}^{2})\over\rho^{2}(\rho^{2}-\rho_{-}^{2})}},\cr K&=\sqrt{{4\rho^{2}(\rho^{2}-\rho_{-}^{2})\over(\rho_{+}^{2}-\rho_{-}^{2})}},\cr Z&=-{\rho_{+}\rho_{-}\over l\rho^{2}}.}$ (4.29)4.29( 4.29 )
The quasi-local angular momentum is given by
$\eqalign{j&={LZ^{\prime}K^{3}\over N},\cr&={2\rho_{+}\rho_{-}\over l}.}$ (4.30)4.30( 4.30 )
Thus, the (Komar) angular momentum is
$J={2\rho_{+}\rho_{-}\over l}.$ (4.31)4.31( 4.31 )
The quasi-local energy is given by
$\eqalign{E&=2(\varepsilon-LK^{\prime}),\cr&=2\varepsilon-{\sqrt{\rho_{+}^{2}-\rho_{-}^{2}}\over l}\left({2\rho^{2}-\rho_{-}^{2}\over\rho^{2}-\rho_{-}^{2}}\right)\sqrt{1-{\rho_{+}^{2}\over\rho^{2}}}.}$ (4.32)4.32( 4.32 )
A particular solution to the Einstein field equations is
$\eqalign{d\widetilde{s}^{2}&={16l^{2}\over(\rho_{+}^{2}-\rho_{-}^{2})^{2}}\rho^{2}d\rho^{2}-{4\over l^{2}(\rho_{+}^{2}-\rho_{-}^{2})}\rho^{4}dt^{2}+{4\over(\rho_{+}^{2}-\rho_{-}^{2})}\rho^{4}dx^{2},\cr e^{-2\phi}&={4\over(\rho_{+}^{2}-\rho_{-}^{2})}\rho^{2},\cr\Lambda&=-{2\over l^{2}}.}$ (4.33)4.33( 4.33 )
We find
$\varepsilon=LK^{\prime}={\sqrt{\rho_{+}^{2}-\rho_{-}^{2}}\over l}.$ (4.34)4.34( 4.34 )
Thus, the quasi-local energy is
$E={\sqrt{\rho_{+}^{2}-\rho_{-}^{2}}\over l}\left[2-\left({2\rho^{2}-\rho_{-}^{2}\over\rho^{2}-\rho_{-}^{2}}\right)\sqrt{1-{\rho_{+}^{2}\over\rho^{2}}}\right],\quad\rho>\rho_{+}.$ (4.35)4.35( 4.35 )
The quasi-local mass is given by
$\eqalign{m&=NE-jZ,\cr&={2\rho_{-}^{2}\over l^{2}}-{4(\rho^{2}-\rho_{+}^{2})\over l^{2}}+{4(\rho^{2}-\rho_{-}^{2})\sqrt{\rho^{2}-\rho_{+}^{2}}\over l^{2}\rho^{2}}.}$ (4.36)4.36( 4.36 )
The (analogous) ADM mass is then given by
$M=m(\rho\to\infty)={2(\rho_{+}^{2}-\rho_{-}^{2})\over l^{2}}.$ (4.37)4.37( 4.37 )
Similarly, for the black hole solutions (4.16) we find that the angular momenta are given by21 Here we set $g_{s}=1$.
$J=2\left({1-\gamma^{2}\over l}\right)=Q,\quad-1\leq\gamma\leq 0.$ (4.38)4.38( 4.38 )
A particular solution to the Einstein field equations is
$\eqalign{d\widetilde{s}^{2}&={(\rho_{+}^{2}-\rho_{-}^{2})\rho^{4}\over 4\gamma l^{2}}dt^{2}-\left({4\gamma\rho^{4}\over\rho_{+}^{2}-\rho_{-}^{2}}\right)\left(dx-{(\rho_{+}^{2}-\rho_{-}^{2})(1+2\lambda\gamma)\over 4l\gamma}dt\right)^{2}+l^{2}\rho^{2}d\rho^{2},\cr e^{-2\phi}&=\rho^{2},\cr\Lambda&=-{2\over l^{2}}.}$ (4.39)4.39( 4.39 )
It gives (assuming $\gamma\neq 0$)
$\varepsilon=LK^{\prime}={4\over l}\sqrt{{-\gamma\over\rho_{+}^{2}-\rho_{-}^{2}}}.$ (4.40)4.40( 4.40 )
The (analogous) ADM masses are given by
$\eqalign{M&=2\left({\rho^{2}_{+}-\rho^{2}_{-}\over l^{2}}\right)\left[1+{(1-\gamma^{2})(1+2\gamma\lambda)\over 4\gamma}\right],\quad-1\leq\gamma<0,\cr&=2\left({\rho^{2}_{+}-\rho^{2}_{-}\over l^{2}}\right)+{2(1-\gamma^{2})\over l}{\widetilde{B}}_{tx}(\rho=\infty),\quad-1\leq\gamma\leq 0,}$ (4.41)4.41( 4.41 )
where
${\widetilde{B}}_{tx}(\rho\to\infty)=\left\{\eqalign{{(\rho_{+}^{2}-\rho_{-}^{2})(1+2\lambda\gamma)\over 4\gamma l}+\left({\rho_{+}^{2}-\rho_{-}^{2}\over 4\rho}\right)^{2}\left({1-\gamma^{2}\over l\gamma^{2}}\right)+{\cal O}(1/\rho^{4}),\quad\gamma\neq 0\cr-{\rho^{2}\over l}+{\cal O}(\rho^{0}),\quad\gamma=0.}\right.$ (4.42)4.42( 4.42 )
Thus, we must choose $-1/2\gamma\leq\lambda<-1/\gamma$ for the (analogous) ADM mass to be positive. Therefore, we must either require the Kalb-Ramond field in (3.15) to go to zero at the boundary or else we have to allow black hole solutions with negative (analogous) ADM masses.
At $\lambda=-1/2\gamma$ we note that the Kalb-Ramond field is zero at infinity. The ADM mass and angular momentum are
$M=2\left({\rho_{+}^{2}-\rho_{-}^{2}\over l^{2}}\right),\quad J=2\left({1-\gamma^{2}\over l}\right)=Q.$ (4.43)4.43( 4.43 )
In the regime $0<\lambda<-1/2\gamma$ the second term in the square bracket (4.41) is negative and thus allows having negative ADM masses. In the cases in which
$00<\lambda<-{1\over 2\gamma}\left(1+{4\gamma\over 1-\gamma^{2}}\right),\quad 2-\sqrt{5}<\gamma<$ 0 . (4.44)4.44( 4.44 )
the corresponding black hole solutions have negative masses. Note this implies for $-1\leq\gamma\leq 2-\sqrt{5}$ the mass is positive for any allowed value of $\lambda$.
A solution with a non-zero asymptotic Kalb-Ramond field is thus equivalent to a solution with no Kalb-Ramond field and no vanishing rotation at asymptotic infinity. In particular, there is dragging, usually known as the Lense-Thirring effect, at infinity, i.e., $Z(r=\infty)\neq 0$. Thus, locally zero angular momentum observers (ZAMOs) [[36]36,,  [37]37] located at asymptotic infinity rotate with a non-zero constant angular velocity $\Omega=Z(r=\infty)$.
Although non-zero $\Omega$ can certainly lower (and raise) the mass, see e.g. the general expression (C.27) of the mass in the appendix, it does not always result in negative mass solutions. A simple example is the BTZ solution with non-vanishing rotation at asymptotic infinity. To allow rotation at infinity we shift $Z$ in (4.29) by a non-zero constant. We denote the constant by $\overline{\lambda}$ and make the change
$Z\rightarrow Z+{\overline{\lambda}\over l}.$ (4.45)4.45( 4.45 )
For the solution to describe a black hole we restrict $\overline{\lambda}$ to the region
$|\overline{\lambda}|<1.$ (4.46)4.46( 4.46 )
The ADM mass is given by
$M={1\over l^{2}}(\rho_{+}^{2}+\rho_{-}^{2}-2\overline{\lambda}\rho_{+}\rho_{-}),\quad-1<\overline{\lambda}<1.$ (4.47)4.47( 4.47 )
Thus, we note that the mass is always positive for all admissible values of $\overline{\lambda}$. The angular momentum does not change and it is still given by (4.24).
The BTZ solution with non-zero rotation $\overline{\lambda}$ at infinity is (or should be viewed as) the usual BTZ solution in a constant Kalb-Ramond field background.
5. Discussion
In this paper we obtained new black hole and black string solutions in the string background ${\cal A}_{3}$ described by the metric $g_{\mu\nu}$, Kalb-Ramond field $B_{\mu\nu}$ and dilaton $\Phi$,
$\eqalign{ds^{2}&=g_{\mu\nu}dx^{\mu}dx^{\nu}=l^{2}(d\theta^{2}-e^{2\phi}{\tenrm cosh}^{2}\theta d\varphi^{2}+e^{2\phi}{\tenrm sinh}^{2}\theta d\psi^{2}),\cr B&=B_{01}d\varphi\wedge d\psi,\quad B_{01}=-{1\over 2}l^{2}e^{2\phi}(\gamma-{\tenrm cosh}(2\theta)),\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|,}$ (5.1)5.1( 5.1 )
where
$e^{-2\phi}=1+\gamma^{2}-2\gamma{\tenrm cosh}(2\theta).$ (5.2)5.2( 5.2 )
The coupling $\gamma$ takes its value in the range $[-1,1]$. The spacetime ${\cal A}_{3}$ (5.1) interpolates between $AdS_{3}$ and (asymptotically) flat spacetime ${\tenrm I\kern-1.79993ptR}\times S^{1}\times{\tenrm I\kern-1.79993ptR}$. It is obtained by applying a sequence of $\alpha^{\prime}$ exact transformations on $AdS_{3}$ [1]. In [1] the deformation is interpreted in general as moving holographic boundaries. The spacetime ${\cal A}_{3}$ is free of curvature singularity when $\gamma$ is negative. Thus, our discussion has been restricted mainly to the case $-1\leq\gamma\leq 0$. For positive $\gamma$ the deformation generates a curvature singularity from the onset [1].
The black string solutions are described by (3.15) which we again write below for convenience. They are obtained from ${\cal A}_{3}$ by applying the simple coordinate transformation (3.13) (and making the change of variable (3.14)). They carry the axion charge per unit length $Q$ (3.19). The black string solutions are
$\eqalign{ds^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-e^{2\phi}{(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})\over l^{2}\rho^{2}}dt^{2}+e^{2\phi}\rho^{2}\left(dx-{\rho_{+}\rho_{-}\over l\rho^{2}}dt\right)^{2},\cr B&=B_{tx}dt\wedge dx,\cr B_{tx}&={1\over 2l}e^{2\phi}\left[\rho_{+}^{2}(1+\lambda+\gamma(1+2\lambda)+\lambda\gamma^{2})+\rho_{-}^{2}(1-\lambda-\gamma(1-2\lambda)-\lambda\gamma^{2})-2\rho^{2}(1+2\lambda\gamma)\right],\cr Q&=e^{-2\Phi}{\star H}=2\left|{1-\gamma^{2}\over lg_{s}^{2}}\right|,\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|,}$ (5.3)5.3( 5.3 )
where
$\eqalign{e^{-2\phi}&={\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\over\rho_{+}^{2}-\rho_{-}^{2}},}$ (5.4)5.4( 5.4 )
$\lambda$ is an arbitrary constant which determines the value of the $B$ field, i.e. $B_{tx}$, at $\rho=\infty$. It does not enter in $H$. However, it has a non-trivial implication. The black string solution (5.3) has inner and outer horizons. This is also a general feature of rotating black holes. The inner horizon is at $\rho=\rho_{-}$ and the outer horizon is at $\rho=\rho_{+}$. In this paper, we in general assume, unless stated otherwise, $x$ is compact and $\rho$ is positive. In the case there is no curvature singularity we identify $x$ so that there is no conical singularity. Its period does not depend on the gauge parameter $\lambda$. To the best of my knowledge the black string solutions (5.3) have not been reported elsewhere in the literature.
For positive $\gamma$, i.e., $0<\gamma\leq 1$, the solution has a curvature singularity at $\rho=\rho_{0}\geq\rho_{+}$ (4.1). Therefore, it cannot be continuously contracted or deformed to the origin without crossing the horizon(s). For $\gamma=1$ the outer horizon becomes singular, see (3.33). The singularity is behind the horizon in the interior region, i.e., $\rho_{0}>\rho_{-}$. Note the interior region is defined in the usual way as the part of the spacetime that contains the singularity. See the Penrose diagrams in Fig. 4 and 5.
For negative $\gamma$, i.e., $-1\leq\gamma<0$, the solution has no singularity if the condition (3.20),
${\rho_{-}\over\rho_{+}}<{1+\gamma\over 1-\gamma},$ (5.5)5.5( 5.5 )
is satisfied. If the above condition is not met, then the solution has a singularity at $\rho=\rho_{0}\leq\rho_{-}$ (4.1). At $\gamma=-1$, (5.3) reduces to the black hole solution (3.25)
$\eqalign{ds^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{1\over 4}{(\rho^{2}-\rho_{+}^{2})(\rho_{+}^{2}-\rho_{-}^{2})\over l^{2}\rho^{2}}dt^{2}+{1\over 4}{\rho^{2}(\rho_{+}^{2}-\rho_{-}^{2})\over\rho^{2}-\rho_{-}^{2}}\left(dx-{\rho_{+}\rho_{-}\over l\rho^{2}}dt\right)^{2},\cr e^{2\Phi}&={g_{s}^{2}\over 4}\left|{\rho_{+}^{2}-\rho_{-}^{2}\over\rho^{2}-\rho_{-}^{2}}\right|.}$ (5.6)5.6( 5.6 )
It has a curvature singularity at $\rho=\rho_{-}$ and an event horizon at $\rho=\rho_{+}$. Thus, the singularity is enclosed inside the horizon. Its (analogous) ADM mass and angular momentum are given by (4.31), (4.37),
$M={2(\rho_{+}^{2}-\rho_{-}^{2})\over l^{2}},\quad J={2\rho_{+}\rho_{-}\over l}.$ (5.7)5.7( 5.7 )
See the Penrose diagrams in Fig. 1, 2 and 3.
It follows from the cases $\gamma=\pm 1$ that both inner and outer horizons can be unstable upon coupling rotating black holes to matter. See below also for a novel case in which a ring singularity forms in between the inner and outer horizons.
In section three, we Lorentz boosted the black string solution (5.3) along the $x$ direction with a specific rapidity (3.36),
$\tanh(2\alpha)={2\rho_{+}\rho_{-}\over\rho_{+}^{2}+\rho_{-}^{2}}.$ (5.8)5.8( 5.8 )
The Kalb-Ramond field is invariant under the Lorentz boost. The resulting metric however now has no cross term and thus static, see (3.37). We subsequently applied T-duality transformation along the $\widehat{x}$ direction. This resulted in the black hole solution (3.44),
$\eqalign{d{\widetilde{s}}^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{+}^{2})\over l^{2}\left[\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\right]}d{\widehat{t}}^{2}\cr&+\left[{\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\over(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{-}^{2})}\right]\cr&\cdot\left\{d{\widetilde{x}}-{(\rho_{+}^{2}-\rho_{-}^{2})\left[\rho_{+}^{2}(1+\gamma)(1+\lambda(1+\gamma))+\rho_{-}^{2}(1-\gamma)(1-\lambda(1-\gamma))-2\rho^{2}(1+2\lambda\gamma)\right]\over 2l\left[\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\right]}d{\widehat{t}}\right\}^{2},\cr e^{-2\widetilde{\Phi}}&=|g_{s}^{-2}(\rho^{2}-\rho_{-}^{2})|,\quad-1\leq\gamma\leq 0,\quad 0<\lambda<-1/\gamma.}$ (5.9)5.9( 5.9 )
To the best of my knowledge this black hole solution has not been reported elsewhere in the literature. The black hole solution has a singularity at $\rho=\rho_{-}$ and an event horizon at $\rho=\rho_{+}$. See the Penrose diagram in Fig. 3. Its (analogous) ADM mass and angular momentum are given by (4.38), (4.41)
$\eqalign{M&=2\left({\rho^{2}_{+}-\rho^{2}_{-}\over l^{2}}\right)\left[1+{(1-\gamma^{2})(1+2\gamma\lambda)\over 4\gamma}\right],\quad-1\leq\gamma<0,\cr&=2\left({\rho^{2}_{+}-\rho^{2}_{-}\over l^{2}}\right)+Q\cdot\Omega,\quad-1\leq\gamma\leq 0,\cr J&=2\left({1-\gamma^{2}\over l}\right)=Q,\quad-1\leq\gamma\leq 0,}$ (5.10)5.10( 5.10 )
where $\Omega$ is the angular velocity at infinity (4.42)
$\Omega=\widetilde{B}_{tx}(\rho=\infty).$ (5.11)5.11( 5.11 )
We note that the black hole depends on the gauge parameter $\lambda$. In particular we can have black holes with negative masses (4.44). The mass is positive if we choose $-1/2\gamma\leq\lambda<-1/\gamma$. Note in the case of BTZ solutions, those solutions with negative masses do not correspond to black holes, see (2.13) and (2.14). We also note that the deformation slows down the rotation [1]. That is, the norm of $J$ decreases as we increase the norm of $\gamma$. The axion charge (per unit length) $Q$ of the black string (with $g_{s}=1$) equals the angular momentum $J$ of the black hole. This is consistent with the work [22].
At $\gamma=-1$ we note that $J=0$ and the mass $M$ is independent of the gauge parameter $\lambda$. $M$ is determined by the radii $\rho_{+}$ and $\rho_{-}$. At $\lambda=-1/2\gamma$, the angular velocity at infinity is zero, i.e., $\Omega=0$. In this case the ADM mass and angular momentum reduce to (4.43)
$M=2\left({\rho_{+}^{2}-\rho_{-}^{2}\over l^{2}}\right),\quad J=2\left({1-\gamma^{2}\over l}\right)=Q.$ (5.12)5.12( 5.12 )
In this paper we also noted that a solution with a non-zero asymptotic Kalb-Ramond field is equivalent to a solution with no Kalb-Ramond field and no vanishing rotation at asymptotic infinity. In general, non-zero rotation allowes negative mass solutions.
In an upcoming work we study the thermodynamics and partition functions of the black hole and black string solutions. We also study the casual structure of the solutions in detail.
In this paper we mainly presented black hole and black string solutions. In a future work we hope to study in greater detail the marginal operator that generates the deformation on the worldsheet.
We also hope to study other string solutions by gluing those solutions of positive $\gamma$ with negative $\gamma$. We give below an example which is obtained by gluing (3.25) and (3.33). It is convenient to introduce first the parameters $0<{\overline{\rho}}_{-}<{\overline{\rho}}_{+}=\rho_{-}<\rho_{+}$. We define $\delta=(\rho_{+}-{\overline{\rho}}_{-})/\rho_{-}$. We assume $\rho_{+}<\sqrt{2}\rho_{-}$. We furthermore assume ${\overline{\rho}}_{-}=\sqrt{2\rho_{-}^{2}-\rho_{+}^{2}}$.
For $\rho\leq{\overline{\rho}}_{+}=\rho_{-}$, the solution is given by
$\eqalign{ds^{2}=&{l^{2}\rho^{2}d\rho^{2}\over({\overline{\rho}}_{+}^{2}-\rho^{2})({\overline{\rho}}_{-}^{2}-\rho^{2})}-{({\overline{\rho}}_{+}^{2}-{\overline{\rho}}_{-}^{2})({\overline{\rho}}_{-}^{2}-\rho^{2})dt^{2}\over 4l^{2}\rho^{2}}+{({\overline{\rho}}_{+}^{2}-{\overline{\rho}}_{-}^{2})\rho^{2}\over 4({\overline{\rho}}_{+}^{2}-\rho^{2})}\left[dx-{1\over l}\left(\delta+{{\overline{\rho}_{+}}{\overline{\rho}_{-}}\over\rho^{2}}\right)dt\right]^{2},\cr B_{tx}&={{\overline{\rho}}_{+}^{2}-{\overline{\rho}}_{-}^{2}\over 4l}={\tenrm const.},\quad Q=0,\cr e^{2\Phi}&={g_{s}^{2}\over 4}\left({\overline{\rho}}_{+}^{2}-{\overline{\rho}}_{-}^{2}\over{\overline{\rho}}_{+}^{2}-\rho^{2}\right).}$ (5.13)5.13( 5.13 )
For $\rho\geq{\overline{\rho}}_{+}=\rho_{-}$, the solution is given by
$\eqalign{ds^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{1\over 4}{(\rho^{2}-\rho_{+}^{2})(\rho_{+}^{2}-\rho_{-}^{2})\over l^{2}\rho^{2}}dt^{2}+{1\over 4}{\rho^{2}(\rho_{+}^{2}-\rho_{-}^{2})\over\rho^{2}-\rho_{-}^{2}}\left(dx-{\rho_{+}\rho_{-}\over l\rho^{2}}dt\right)^{2},\cr B_{tx}&=-\left({\rho_{+}^{2}-\rho_{-}^{2}\over 4l}\right)={\tenrm const.},\quad Q=0,\cr e^{2\Phi}&={g_{s}^{2}\over 4}\left(\rho_{+}^{2}-\rho_{-}^{2}\over\rho^{2}-\rho_{-}^{2}\right).}$ (5.14)5.14( 5.14 )
The scalar curvature of the solution (which is described by (5.13) and (5.14)) is $R=-4\xi^{2}\cdot(\rho_{+}^{2}/l^{2})\cdot(1/|\rho^{2}-\rho_{-}^{2}|)$, where $\xi^{2}=1-(\rho_{-}/\rho_{+})^{2}$. The Kretschmann scalar is $K=R^{2}$. Thus, the solution has the peculiar feature that it contains a ring curvature singularity at $\rho={\overline{\rho}}_{+}=\rho_{-}$ in between the inner and outer horizons. The inner horizon is at $\rho={\overline{\rho}}_{-}$. The outer horizon is at $\rho=\rho_{+}$. Note they are not related by sign reflection, i.e. ${\overline{\rho}}_{-}\neq-\rho_{+}$. The Kalb-Ramond field is a constant (field). It can be fixed to zero. The Penrose diagram of the solution is given in Fig. 6. It is obtained by combining Fig. 3 and 5.
Fig. 6: The plot depicts the global structure of the string solution described by (5.13) and (5.14). The solution has a singularity at $\rho=\rho_{-}={\overline{\rho}}_{+}$. It is represented by the zigzag lines.
Note usually the diagram ends at $\rho={\overline{\rho}}_{+}=\rho_{-}$, however here, since there are no closed time-like curves, the diagram extends outward to infinity and inward to zero. Therefore, the maximal extension simply extends infinitely in both the vertical directions.
We expect the black hole solution (described by (5.13) and (5.14)) to have a generalization in (3 + 1) dimensions with similar feature. That is, with a curvature singularity in between the inner and outer horizons. We hope to study the corrections in $g_{s}$ and their roles in resolving the singularity in a future work.
Acknowledgements: This work is supported by the Department of Atomic Energy under project no. RTI4001.
Appendix A. Invariant metric on $AdS_{3}$
The (invariant) metric of the Lie algebra $su(1,1)\cong sl(2,{\tenrm I\kern-1.79993ptR})$ is given by the Killing form $ds^{2}={\tenrm tr}\left(g^{-1}dgg^{-1}dg\right)$ (up to a constant factor), where $g^{-1}dg$ is the (left invariant) Maurer-Cartan one form and $g$ is $SU(1,1)\cong SL(2,R)$ group element.
A convenient choice of basis for $su(1,1)$ is $a_{1}=-{1\over 2}\sigma_{1},a_{2}=-{1\over 2}\sigma_{2},a_{3}={i\over 2}\sigma_{3}$, where $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$ are the Pauli spin matrices. They are given by
[…]
We parametrize (using the Cartan decomposition) the group element $g\in SU(1,1)$ as
$\eqalign{g(\theta,\varphi,\psi)&=e^{\left(\varphi+\psi\right)a_{3}}e^{2\theta a_{1}}e^{\left(\varphi-\psi\right)a_{3}},\cr&=e^{i\left(\varphi+\psi\over 2\right)\sigma_{3}}e^{-\theta\sigma_{1}}e^{i\left(\varphi-\psi\over 2\right)\sigma_{3}},\cr&=\pmatrix{e^{i\varphi}{\tenrm cosh}\theta&-e^{i\psi}{\tenrm sinh}\theta\cr-e^{-i\psi}{\tenrm sinh}\theta&e^{-i\varphi}{\tenrm cosh}\theta},}$ (A.2)A.2( A. 2 )
here $0\leq\theta<\infty,0\leq\varphi<2\pi,0\leq\psi<2\pi$. This choice gives the invariant metric
$ds^{2}=d\theta^{2}-{\tenrm cosh}^{2}\theta d\varphi^{2}+{\tenrm sinh}^{2}\theta d\psi^{2}.$ (A.3)A.3( A. 3 )
On the universal cover $\varphi\in{\tenrm I\kern-1.79993ptR}$.
However, as we see in section two, it is the real form of the Lie algebra $su(1,1)$ (i.e., now the field is ${\tenrm I\kern-1.79993ptR}$ however the basis can be complex) or simply the lie algebra $sl(2,R)$ which is in particular relevant for our discussion in relation to the exterior region of the BTZ black hole. Therefore, we obtain below the metric using $sl(2,{\tenrm I\kern-1.79993ptR})$. A convenient choice of basis is $b_{1}=-{1\over 2}\sigma_{1},b_{2}={i\over 2}\sigma_{2},b_{3}=-{1\over 2}\sigma_{3}$. The analogous parametrization is
$\eqalign{g(\overline{\theta},\overline{\varphi},\overline{\psi})&=e^{\left(\overline{\varphi}+\overline{\psi}\right)b_{3}}e^{2\overline{\theta}b_{1}}e^{\left(\overline{\varphi}-\overline{\psi}\right)b_{3}},\cr&=e^{-\left(\overline{\varphi}+\overline{\psi}\over 2\right)\sigma_{3}}e^{-\overline{\theta}\sigma_{1}}e^{-\left(\overline{\varphi}-\overline{\psi}\over 2\right)\sigma_{3}},\cr&=\pmatrix{e^{-\overline{\varphi}}{\tenrm cosh}\overline{\theta}&-e^{-\overline{\psi}}{\tenrm sinh}\overline{\theta}\cr-e^{\overline{\psi}}{\tenrm sinh}\overline{\theta}&e^{\overline{\varphi}}{\tenrm cosh}\overline{\theta}},}$ (A.4)A.4( A. 4 )
here $\overline{\theta}\in{\tenrm I\kern-1.79993ptR},\overline{\varphi}\in{\tenrm I\kern-1.79993ptR},\overline{\psi}\in{\tenrm I\kern-1.79993ptR}$. In this choice the (invariant) metric is given by
$d{\overline{s}}^{2}=d\overline{\theta}^{2}+{\tenrm cosh}^{2}\overline{\theta}d\overline{\varphi}^{2}-{\tenrm sinh}^{2}\overline{\theta}d\overline{\psi}^{2}.$ (A.5)A.5( A. 5 )
The metric (A.5) can be equivalently obtained from (A.3) formally by the analytic continuation
$\varphi=i\overline{\varphi},\quad\psi=i\overline{\psi}.$ (A.6)A.6( A. 6 )
We simply replace $\theta$ with $\overline{\theta}\in{\tenrm I\kern-1.79993ptR}$.
Appendix B. Bosonic T-duality
We briefly review T-duality in Type II bosonic string theory. See [[38]38[39]--[40]40,,  [41]41] for detailed discussions. See also Appendix A in [1].
Consider the world-sheet sigma model action
$S:=\int d^{2}z{\cal L},\quad{\cal L}=\partial X^{m}\Sigma_{mn}(X)\overline{\partial}X^{n},\quad\Sigma_{mn}:=G_{mn}+B_{mn},$ (B.1)B.1( B. 1 )
where $G_{mn}=G_{nm}$ are the components of the metric $G$ and $B_{mn}=-B_{mn}$ are the components of the Kalb-Ramond two-form $B$. $z=\tau+\sigma$ and ${\overline{z}}=\tau-\sigma$ are the world-sheet coordinates. We omitted a term linear in the dilaton field. The dilaton $\Phi$ enters at the quantum level [[42]42,,[43]43].
We assume the background fields $G_{mn}$, $B_{mn}$ and $\Phi$ are independent of $X^{1}$. Thus, the transformation
$X^{1}\to X^{1}+c,$ (B.2)B.2( B. 2 )
is a symmetry of the action, where $c$ is a constant. Therefore, we can couple the theory to a gauge field $A_{a}$ on the worldsheet. The gauged Lagrangian density is
${\cal L}=\partial X^{m}\Sigma_{mn}\overline{\partial}X^{n}+\partial X^{m}\Sigma_{m1}{\overline{A}}+A\Sigma_{1n}\overline{\partial}X^{n}+A\Sigma_{11}{\overline{A}}+\chi(\partial{\overline{A}}-\overline{\partial}A),$ (B.3)B.3( B. 3 )
where the field $\chi$ is a Lagrange multiplier.
We assume $X^{1}$ is periodic and compact, i.e., at fixed worldsheet time it parametrizes a circle. Thus, the symmetry (B.2) is a global $U(1)$ (zero-form) symmetry. We fix the gauge by assuming $X^{1}={\tenrm const.}$. This gives
$S=\int d^{2}z{\cal L},\quad{\cal L}=\partial X^{\widehat{m}}\Sigma_{{\widehat{m}}{\widehat{n}}}\overline{\partial}X^{\widehat{n}}+\partial X^{\widehat{m}}\Sigma_{{\widehat{m}}1}{\overline{A}}+A\Sigma_{1{\widehat{n}}}\overline{\partial}X^{\widehat{n}}+A\Sigma_{11}{\overline{A}}+\chi(\partial{\overline{A}}-\overline{\partial}A),$ (B.4)B.4( B. 4 )
where $\widehat{m}$ takes all values of $m$ except $m=1$. Integrating out the field $\chi$, i.e., eliminating it using its equation of motion, gives back (B.1). In general a one form on $S^{1}$ is not exact. However, since $X^{1}$ is compact and thus it winds around in either directions integer times, we can still write the gauge potential on $S^{1}$ as the gradient of a compact scalar $A=\partial\chi^{1},{\overline{A}}={\overline{\partial}}\chi^{1}$ and make the identification $\chi^{1}\sim X^{1}$. This will get us back to the original action.
If we instead perform integration by parts, it leads to
$S^{\prime}=\int d^{2}z{\cal L}^{\prime},\quad{\cal L}^{\prime}=\partial X^{\widehat{m}}\Sigma_{{\widehat{m}}{\widehat{n}}}\overline{\partial}X^{\widehat{n}}+A\Sigma_{11}{\overline{A}}-\left(\partial\chi-\partial X^{\widehat{m}}\Sigma_{{\widehat{m}}1}\right){\overline{A}}+A\left(\Sigma_{1{\widehat{n}}}\overline{\partial}X^{\widehat{n}}+\overline{\partial}\chi\right).$ (B.5)B.5( B. 5 )
Integrating out the gauge fields $A$ and $\overline{A}$ using their equations of motion gives the T-dual Lagrangian density
${\cal L}^{\prime}=\partial X^{\widehat{m}}\sigma_{{\widehat{m}}{\widehat{n}}}\overline{\partial}X^{\widehat{n}}+\partial\chi\sigma_{\chi{\widehat{m}}}{\overline{\partial}X}^{\widehat{m}}+\partial X^{\widehat{m}}\sigma_{{\widehat{m}}\chi}\overline{\partial}\chi+\partial\chi\sigma_{\chi\chi}\overline{\partial}\chi,$ (B.6)B.6( B. 6 )
where
$\sigma_{\chi\chi}={1\over\Sigma_{11}},\quad\sigma_{\chi\widehat{m}}={\Sigma_{1\widehat{m}}\over\Sigma_{11}},\quad\sigma_{{\widehat{n}}\chi}=-{\Sigma_{{\widehat{n}}1}\over\Sigma_{11}},\quad\sigma_{{\widehat{m}}{\widehat{n}}}=\Sigma_{{\widehat{m}}{\widehat{n}}}-{\Sigma_{{\widehat{m}}1}\Sigma_{1{\widehat{n}}}\over\Sigma_{11}}.$ (B.7)B.7( B. 7 )
The Jacobian factor introduced in the path integral measure by the change of variables leads to a shift of the dilaton. The new dilaton $\phi$ is
$\phi=\Phi-{1\over 2}\log\left(\Sigma_{11}\right),\quad\Sigma_{11}=G_{11}.$ (B.8)B.8( B. 8 )
The new background fields are arranged as
$\sigma_{{\widehat{m}}{\widehat{n}}}=g_{{\widehat{m}}{\widehat{n}}}+b_{{\widehat{m}}{\widehat{n}}},\quad g_{{\widehat{m}}{\widehat{n}}}=g_{{\widehat{n}}{\widehat{m}}},\quad b_{{\widehat{m}}{\widehat{n}}}=-b_{{\widehat{n}}{\widehat{m}}},$ (B.9)B.9( B. 9 )
and similar expressions for $\sigma_{\chi{\widehat{n}}},\ \sigma_{{\widehat{n}}\chi},\ \sigma_{\chi\chi}$. Thus, the components of the new metric $g$ and two-form $b$ field are given in terms of $G_{mn}$ and $B_{mn}$ by the relations
$g_{\chi\chi}={1\over G_{11}},\quad g_{\chi{\widehat{m}}}={B_{1{\widehat{m}}}\over G_{11}},\quad g_{{\widehat{n}}\chi}=-{B_{{\widehat{n}}1}\over G_{11}},\quad g_{{\widehat{m}}{\widehat{n}}}=G_{{\widehat{m}}{\widehat{n}}}-{1\over G_{11}}\left(G_{{\widehat{m}}1}G_{1{\widehat{n}}}+B_{{\widehat{m}}1}B_{1{\widehat{n}}}\right),$ (B.10)B.10( B. 10 )
and
$b_{\chi{\widehat{m}}}={G_{1{\widehat{m}}}\over G_{11}},\quad b_{{\widehat{n}}\chi}=-{G_{{\widehat{n}}1}\over G_{11}},\quad b_{{\widehat{m}}{\widehat{n}}}=B_{{\widehat{m}}{\widehat{n}}}-{1\over G_{11}}\left(G_{{\widehat{m}}1}B_{1{\widehat{n}}}+B_{{\widehat{m}}1}G_{1{\widehat{n}}}\right).$ (B.11)B.11( B. 11 )
The relations (B.8), (B.10) and (B.11) are also known as the Buscher rules. From the worldsheet perspective, T-dulity is an exact equivalence of CFTs.
Appendix C. The Brown-York quasi-local energy and mass
In this appendix we review the Brown-York quasi-local energy and mass [5]. We primarily follow the work [5], but see also [4] for a related insightful discussion. We specialize the discussion for definiteness to $(2+1)$ dimensions. We furthermore restrict ourselves to axially symmetric stationary solutions. Therefore, we can directly use the formulae in section four of the paper.
The Brown-York definition of energy and mass is based on the Hamilton-Jacobi formulation of general relativity. It requires foliating the spacetime into a one-parameter family of space-like hyper-surfaces and thus defining a time direction. It also requires adding the Gibbons-Hawking-York (GHY) boundary term to the usual action for the gravitational field [[44]44,,[45]45]. The boundary term ensures that only the metric is fixed at the boundary. It avoids the need to fix the first derivatives of the metric.
The general metric (relevant for our purpose) is of the form
$ds^{2}=-A^{2}dt^{2}+{1\over B^{2}}dr^{2}+C^{2}(d\theta+Zdt)^{2},$ (C.1)C.1( C. 1 )
where $A,B,C$ and $Z$ are functions of only the radial coordinate $r$. The coordinate $\theta$ is an angle variable and it has $2\pi$ periodicity. The spacetime contains no inner boundaries. It can have event horizons [4].22 In the case the spacetime contains other null boundaries we include appropriate terms in the action functional that contribute on these boundaries. This will not concern us in this paper. We refer to [46] and references therein for such cases. We assume the metric is minimally coupled to a scalar. Therefore, the scalar action contains no derivatives of the metric. We also assume the scalar depends only on $r$.
Let $u^{\mu}$ denote the future pointing time-like unit normal to the $t={\tenrm constant}$ two-space denoted by $\Sigma_{t}$. $t$ labels the two-spaces $\Sigma_{t}$ and it defines the time direction or coordinate. This particular slicing is the usual ADM slicing of spacetime [47]. We denote the (intrinsic) metric on $\Sigma_{t}$ by $h_{ij}$.23 For convenience, here and below we use lower latin letters $i=t,r,\theta$ to denote the intrinsic metrics and we use greek letters $\mu=t,r,\theta$ to denote the induced metrics. We denote the boundary of $\Sigma_{t}$ by $B_{t}$. The world-line or history of $B_{t}$ is denoted ${}^{2}B$. Let $n^{\mu}$ is the outward pointing space-like unit normal to the two-boundary ${}^{2}B$. For simplicity, we assume on the time-like boundary ${}^{2}B$ that $n^{\mu}u_{\mu}|_{{}^{2}B}=0$. That is, the surface $\Sigma_{t}$ meets the boundary ${}^{2}B$ orthogonally. This is not however essential or necessary for the analysis below [[48]48]. This is in part because we are ultimately interested in the case ${}^{2}B$ is at radial infinity where the spacetime (C.1) is (conformally) flat. We denote the metric on ${}^{2}B$ by $\gamma_{ij}$. We denote the metric on $B_{t}$ which is the intersection of ${}^{2}B$ and $\Sigma_{t}$ by $\sigma_{ij}$.
The time-like and space-like normal vector fields are
$\eqalign{u_{\mu}=-A\delta^{t}_{\mu},\quad u^{\mu}=g^{\mu\nu}u_{\nu}={1\over A}\delta^{\mu}_{t}-{Z\over A}\delta^{\mu}_{\theta},\cr n_{\mu}={1\over B}\delta^{r}_{\mu},\quad n^{\mu}=g^{\mu\nu}n_{\nu}=B\delta^{\mu}_{r}.}$ (C.2)C.2( C. 2 )
The induced metrics are
$\eqalign{h_{\mu\nu}&=g_{\mu\nu}+u_{\mu}u_{\nu},\cr&=g_{\mu\nu}+A^{2}\delta^{t}_{\mu}\delta^{t}_{\nu},\cr\gamma_{\mu\nu}&=g_{\mu\nu}-n_{\mu}n_{\nu},\cr&=g_{\mu\nu}-{1\over B^{2}}\delta^{r}_{\mu}\delta^{r}_{\nu},\cr\sigma_{\mu\nu}&=g_{\mu\nu}+u_{\mu}u_{\nu}-n_{\mu}n_{\nu},\cr&=g_{\mu\nu}+A^{2}\delta^{t}_{\mu}\delta^{t}_{\nu}-{1\over B^{2}}\delta^{r}_{\mu}\delta^{r}_{\nu}.}$ (C.3)C.3( C. 3 )
The extrinsic curvature to the surface $\Sigma_{t}$ is
$\eqalign{K_{\mu\nu}&=-h^{\alpha}_{\mu}\nabla_{\alpha}u_{\nu},\cr&={C^{2}Z(\partial_{r}Z)\over 2A}(\delta^{t}_{\nu}\delta^{r}_{\mu}+\delta^{t}_{\mu}\delta^{r}_{\nu})+{1\over 2A}C^{2}(\partial_{r}Z)(\delta^{\theta}_{\nu}\delta^{r}_{\mu}+\delta^{\theta}_{\mu}\delta^{r}_{\nu}),\cr&={C^{2}Z(\partial_{r}Z)\over A}\delta^{t}_{(\nu}\delta^{r}_{\mu)}+{1\over A}C^{2}(\partial_{r}Z)\delta^{\theta}_{(\nu}\delta^{r}_{\mu)}.}$ (C.4)C.4( C. 4 )
The extrinsic curvature to the boundary ${}^{2}B$ is
$\eqalign{\Theta_{\mu\nu}=&-\gamma^{\alpha}_{\mu}\nabla_{\alpha}n_{\nu},\cr=&-BC\left[C(\partial_{r}Z)+2Z(\partial_{r}C)\right]\delta^{t}_{(\mu}\delta^{\theta}_{\nu)}-BC(\partial_{r}C)\delta^{\theta}_{\mu}\delta^{\theta}_{\nu}\cr&+B\left[A(\partial_{r}A)-C^{2}Z(\partial_{r}Z)-Z^{2}C(\partial_{r}C)\right]\delta^{t}_{\mu}\delta^{t}_{\nu}.}$ (C.5)C.5( C. 5 )
The extrinsic curvature to the boundary $B_{t}$ is
$\eqalign{k_{\mu\nu}&=-\sigma^{\alpha}_{\mu}\nabla_{\alpha}n_{\nu},\cr&=-BC(\partial_{r}C)\delta^{\theta}_{\mu}\delta^{\theta}_{\nu}-Z^{2}BC(\partial_{r}C)\delta^{t}_{\mu}\delta^{t}_{\nu}-2ZBC(\partial_{r}C)\delta^{\theta}_{(\mu}\delta^{t}_{\nu)},\cr&=\sigma^{\alpha}_{\mu}\sigma^{\beta}_{\nu}\Theta_{\alpha\beta}.}$ (C.6)C.6( C. 6 )
The trace $k$ is
$k=\sigma^{\mu\nu}k_{\mu\nu}=-BC(\partial_{r}C)\sigma^{\theta\theta}=-{B(\partial_{r}C)\over C}.$ (C.7)C.7( C. 7 )
We denote the canonical momenta on ${}^{2}B$ conjugate to $\gamma_{ij}$ by $\pi^{ij}$. The momenta $\pi^{ij}$ contain no scalar contribution since the scalar field is coupled to gravity minimally. Thus, they are purely gravitational. The gravitational canonical momenta are given by
$\eqalign{\pi^{ij}&=-{1\over 2\kappa^{2}}\sqrt{-\gamma}\left(\Theta\gamma^{ij}-\Theta^{ij}\right),}$ (C.8)C.8( C. 8 )
where $\Theta_{ij}=\gamma^{\mu}_{i}\gamma^{\nu}_{j}\Theta_{\mu\nu}$ and $\Theta=g^{\mu\nu}\Theta_{\mu\nu}=\gamma^{ij}\Theta_{ij}$ is the trace. In general, $\pi^{ij}$ defines a total energy-momentum tensor on the surface ${}^{2}B$. The surface energy-momentum tensor $\tau_{ij}$ is
$\tau^{ij}:={2\over\sqrt{-\gamma}}{\delta S_{\tenrm on-shell}\over\delta\gamma_{ij}}={2\over\sqrt{-\gamma}}\pi^{ij},$ (C.9)C.9( C. 9 )
where $S_{\tenrm on-shell}$ is the on-shell action functional. It satisfies the momentum constraints
${\cal D}_{i}\tau^{ij}=-T^{\mu\nu}n_{\mu}\gamma^{j}_{\nu}\equiv-T^{nj},$ (C.10)C.10( C. 10 )
where $T^{\mu\nu}$ is the matter energy-momentum tensor and ${\cal D}_{i}$ is the covariant derivative compatible with the metric $\gamma_{ij}$, i.e., ${\cal D}_{i}\gamma^{jk}=0$. It is defined by projecting $\nabla_{\mu}$ onto ${}^{2}B$, i.e., ${\cal D}_{i}=\gamma_{i}^{\lambda}\nabla_{\lambda}$. The proper surface energy density $\varepsilon$ is
$\eqalign{\varepsilon&\equiv u_{i}u_{j}\tau^{ij}={1\over\kappa^{2}}k,\cr&={1\over\kappa^{2}}\cdot-{B(\partial_{r}C)\over C}.}$ (C.11)C.11( C. 11 )
The proper surface momentum density $j_{i}$ is
$\eqalign{j_{k}&\equiv-\sigma_{ki}u_{j}\tau^{ij}={1\over\kappa^{2}}\sigma^{i}_{k}n^{j}K_{ij},\cr&={BC^{2}(\partial_{r}Z)\over 2\kappa^{2}A}\left(\delta^{\theta}_{k}+Z\delta^{t}_{k}\right).}$ (C.12)C.12( C. 12 )
For a minimally coupled scalar $\phi$ with action (4.6) the energy momentum tensor is given by (4.12),
$\kappa^{2}T_{\mu\nu}=4\nabla_{\mu}\phi\nabla_{\nu}\phi-2g_{\mu\nu}\left(\nabla\phi\right)^{2}-{1\over 2}g_{\mu\nu}V(\phi).$ (C.13)C.13( C. 13 )
Thus, the projection onto ${}^{2}B$ is
$\eqalign{T^{nj}&=T^{\mu\nu}n_{\mu}\gamma^{j}_{\nu},\cr&={1\over B}\left(T^{rt}\delta^{j}_{t}+T^{r\theta}\delta^{j}_{\theta}\right).}$ (C.14)C.14( C. 14 )
We have (since $\gamma^{j}_{r}=0$)
$\eqalign{T^{rt}&={B^{2}\over N^{2}}\left(-T_{rt}+ZT_{r\theta}\right),\cr T^{r\phi}&={B^{2}\over N^{2}}\left[ZT_{rt}+\left({A^{2}\over C^{2}}-Z^{2}\right)T_{r\theta}\right].}$ (C.15)C.15( C. 15 )
In the case the scalar $\phi$ is a function of only $r$, which is the case in this paper, and the metric is of the form (C.1), we have
$T_{rt}=0,\quad T_{r\theta}=0.$ (C.16)C.16( C. 16 )
Therefore,
$T^{nj}=0.$ (C.17)C.17( C. 17 )
Let $\xi^{i}$ denote a Killing vector field associated with an isometry of the boundary two-metric $\gamma_{ij}$. Thus, it solves
${\cal D}_{(i}\xi_{j)}=0.$ (C.18)C.18( C. 18 )
In the case $T^{ni}=0$ (or $T^{ni}\xi_{i}=0$) on the surface ${}^{2}B$, which is the case in this paper (see (C.17)), it follows from (C.10) that the quantity,
[…]
is conserved, i.e., $Q_{\xi}(B_{t_{0})}-Q_{\xi}(B_{t_{1}})=-\int_{{}^{2}B}d^{2}x\sqrt{-\gamma}T^{ni}\xi_{i}=0$.
We now give the conserved charges in terms of the components of the metric. The energy associated with the Killing field $\xi^{i}=u^{i}=\gamma_{i}^{\mu}u_{\mu}$24 Note ${\cal D}_{(i}\xi_{j)}=\gamma_{(i}^{\lambda}\nabla_{\lambda}\gamma_{j)}^{\mu}u_{\mu}=0$. is given by
[…]
The mass associated with the Killing field $\xi^{i}=\left(\partial_{t}\right)^{i}=\delta^{i}_{t}$ is given by
[…]
Note the mass corresponds as in asymptotically flat spacetimes to the Hamiltonian.
In gravity theories physical quantities such as energy are defined only with respect to a reference. Thus, we must specify a reference spacetime ${\cal M}_{0}(g_{0},\phi_{0};\Lambda)$, where $g_{0}$ denotes the metric and $\phi_{0}$ collectively denotes the matter fields on ${\cal M}_{0}$. $\Lambda$ is possible cosmological constant.
The reference spacetime ${\cal M}_{0}(g_{0},\phi_{0};\Lambda)$ is a particular solution to Einstein equations. It is chosen such that the deviations or fluctuations
$\eqalign{h_{\mu\nu}:=g_{\mu\nu}-(g_{0})_{\mu\nu},\cr\chi:=\phi-\phi_{0},}$ (C.23)C.23( C. 23 )
vanish smoothly at the boundary $B_{t}$ [3]. Note $B_{t}$ is the boundary of the $t={\tenrm constant}$ spacetime slice. In the case ${\cal M}(g,\phi;\Lambda)$ is asymptotically flat spacetime, the reference spacetime ${\cal M}_{0}(g_{0},\phi_{0};\Lambda)$ is usually taken to be the Minkowski spacetime. In this paper in all the cases we study the reference metrics are conformally flat spacetimes.
We take a reference metric $g_{0}$ of the form
$ds^{2}_{0}=-A_{0}^{2}dt^{2}+{1\over B^{2}_{0}}dr^{2}+C^{2}_{0}(d\theta+Z_{0}dt)^{2}.$ (C.24)C.24( C. 24 )
We will assume $A_{0},B_{0},C_{0}$ and $Z_{0}$ are functions of only $r$. The scalar field depends only on $r$ and it is coupled to the metric minimally. At the boundary $B_{t}$, $h_{\mu\nu}=0$ and $\chi=0$ (C.23). That is, $A(r^{\star})=A_{0}(r^{\star}),B(r^{\star})=B_{0}(r^{\star}),C(r^{\star})=C_{0}(r^{\star}),Z(r^{\star})=Z_{0}(r^{\star})$ and $\phi(r^{\star})=\phi_{0}(r^{\star})$, where $r^{\star}$ is the location of $B_{t}$.
The reference metrics $g_{0}$ that we consider in this paper have $Z_{0}={\tenrm constant}$. Accordingly, since $\partial_{r}Z_{0}=0$, the (quasi-local) angular momentum remains unchanged. It is still given by (C.21). The (quasi-local) energy is given by26 In general it is quasi-local in the sense that it is local only on the (hyper-)surface $\Sigma_{t}$, i.e., the $t={\tenrm constant}$ spacetime slice. It follows, in $(2+1)$ dimensions, it is defined as a local quantity.
$E={2\pi\over\kappa^{2}}\left(B_{0}\partial_{r}C_{0}-B\partial_{r}C\right).$ (C.25)C.25( C. 25 )
We implicitly assume the limit $r\to r^{\star}$ exists. The (quasi-local) mass is
$\eqalign{m&=AE-Zj,\cr&={\pi\over\kappa^{2}}\left[2A\left(B_{0}\partial_{r}C_{0}-B\partial_{r}C\right)-{ZBC^{3}\partial_{r}Z\over A}\right].}$ (C.26)C.26( C. 26 )
The angular momentum and (analogous or physical) ADM mass are defined by taking $r$ to infinity or sending $B_{t}$ to infinity [[3]3,,[4]4]
$\eqalign{J=\lim_{r\to\infty}j(r),\cr M=\lim_{r\to\infty}m(r).}$ (C.27)C.27( C. 27 )
In section four of the paper we set $\kappa^{2}=\pi$ (in some length unit) for simplicity. This is equivalent to $18G_{N}=$ 1 (in some length unit).
Appendix D. Kruskal coordinates
The metrics which we study in this paper are in general of the form $ds^{2}=\Delta^{2}d{\widehat{s}}^{2}$, where the factor $\Delta^{2}$ is (manifestly) positive and regular everywhere.27 Except at a curvature singularity. The line element $d{\widehat{s}}^{2}$ is of two types. One type has the form (see, e.g., (3.15) and (4.14))
$d{\widehat{s}}^{2}=\Delta^{-2}f^{-1}d\rho^{2}-fdt^{2}+\rho^{2}(d\phi+gdt)^{2},$ (D.1)D.1( D. 1 )
and the other has (in general with different $f,g$ and $\Delta$) the form (see, e.g., (3.24) and (3.48))
$d{\widehat{s}}^{2}=\Delta^{-2}f^{-1}d\rho^{2}-fdt^{2}+(d\phi+gdt)^{2}.$ (D.2)D.2( D. 2 )
The functions $f,g$ and $\Delta$ depend only on $\rho$. $\phi$ is angular coordinate. The function $f$ has either one or two simple zeros or poles. The scalar curvature is finite at the zero(s). At large $\rho$ both are conformally flat. Thus, each of their Penrose diagrams is similar at large $\rho$ to that of a flat spacetime.
Around each zero we can introduce a coordinate patch to analytically continue the metric. We now show how to continue the metric (D.1) across the zero(s). The analysis is identical for (D.2). We follow [50] and [9].
We show that around a zero of $f$ which we denote by $\rho_{0}$ the metric (D.1) can be put into the form
$d{\widehat{s}}^{2}=\Omega^{2}\left(du^{2}-dv^{2}\right)+\rho^{2}(d\varphi+(g-g_{0})dt)^{2},\quad\varphi=\phi+g_{0}t,$ (D.3)D.3( D. 3 )
where $t=t(u,v)$ and $g_{0}=g(\rho_{0})$. The factor $\Omega^{2}(u,v)$ is regular at the zero. Thus, (D.3) is a continuation of (D.1) across the surface $\rho=\rho_{0}$. In terms of $u(\rho,t)$ and $v(\rho,t)$ light cones at the zero are lines with slope ±$\pm 1$. We show (D.3) by finding $u(t,\rho)$ and $v(t,\rho)$.
By comparing (D.1) and (D.3) we find
$\matrix{\Omega^{2}(u_{\rho}^{2}-v_{\rho}^{2})&=&\Delta^{-2}f^{-1},\cr\Omega^{2}(u_{t}^{2}-v_{t}^{2})&=&-f,\cr u_{\rho}u_{t}-v_{\rho}v_{t}&=&0,}$ (D.4)D.4( D. 4 )
where $u_{\rho}=\partial u/\partial\rho$, $u_{t}=\partial u/\partial t$. Similar notations apply for $v$. We eliminate $\Omega$ and find a relation for $u_{t}(v_{t})$ and $v_{\rho}(u_{\rho})$. We first take a ratio of the first two equations, and then use the last equation to simplify it. The result is
${u_{t}^{2}-v_{t}^{2}\over u_{\rho}^{2}-v_{\rho}^{2}}=-{u_{t}^{2}\over v_{\rho}^{2}}=-f^{2}\Delta^{2}.$ (D.5)D.5( D. 5 )
Thus,
$u_{t}={f\Delta}v_{\rho},\quad v_{t}={f\Delta}u_{\rho}.$ (D.6)D.6( D. 6 )
We now introduce a new radial coordinate $\rho^{\star}$, defined by
$d\rho^{\star}=f^{-1}\Delta^{-1}d\rho.$ (D.7)D.7( D. 7 )
In terms of $t$ and (the tortoise coordinate) $\rho^{\star}$, the equations (D.6) become
$u_{t}=v_{\rho^{\star}},\quad v_{t}=u_{\rho^{\star}}.$ (D.8)D.8( D. 8 )
The general solution of these equations is
$u=h_{+}(\rho^{\star}+t)+h_{-}(\rho^{\star}-t),\quad v=h_{+}(\rho^{\star}+t)-h_{-}(\rho^{\star}-t),$ (D.9)D.9( D. 9 )
where $h_{\pm}$ are arbitrary functions of (the Eddington-Finkelstein coordinates) $\rho^{\star}\pm t$. We take the Kruskal ansatz,
$h_{\pm}=A_{\pm}e^{\kappa(\rho^{\star}\pm t)},$ (D.10)D.10( D. 10 )
where $A_{\pm}$ and $\kappa$ are constants. This gives the factor
$\Omega^{2}=-{f\over u_{t}^{2}-v_{t}^{2}}={f\over 4A_{+}A_{-}\kappa^{2}}e^{-2\kappa\rho^{\star}}.$ (D.11)D.11( D. 11 )
$\kappa$ is chosen such that $\Omega^{2}$ is regular, i.e. non-singular, at the zero of $f$ and on the patch around it. Since the zero or pole is simple, the constant $\kappa$ exists. The Kruskal coordinates are now given by (with $A=A_{-}=A_{+}$)
$u=2Ae^{\kappa\rho^{\star}}{\tenrm cosh}(\kappa t),\quad v=2Ae^{\kappa\rho^{\star}}{\tenrm sinh}(\kappa t).$ (D.12)D.12( D. 12 )
Without loss of generality we can take $A=\pm 1$. The inverse transformations are
$\rho^{\star}={1\over 2\kappa}\ln\left({u^{2}-v^{2}\over 4A^{2}}\right),\quad t={1\over 2\kappa}\tanh^{-1}\left({2uv\over u^{2}+v^{2}}\right).$ (D.13)D.13( D. 13 )
Penrose coordinates $(U,V)$ are given by the usual relations
$u+v=\tan\left({U+V\over 2}\right),\quad u-v=\tan\left({U-V\over 2}\right).$ (D.14)D.14( D. 14 )
In a situation where there are more then one zero, the analysis involves introducing several patches, one for each zero. Overlapping regions are identified. The maximal extension requires including and gluing together infinite copies of the patches.
We now give an example to illustrate the procedure. We consider the case (3.20). We find it convenient to define
$\lambda={\rho_{-}\over\rho_{+}},\quad c_{1}^{2}={l^{2}\over\rho_{+}^{2}(1+\mu)}\sqrt{{\mu^{2}-\lambda^{2}\over 1-\lambda^{2}}},\quad c_{2}^{2}={2(1-\mu^{2})\over\mu(\mu^{2}-\lambda^{2})},\quad x={\rho\over\rho_{+}},\quad x^{\star}={\rho^{\star}\over\rho_{+}},$ (D.15)D.15( D. 15 )
where $\mu=(1+\gamma)/(1-\gamma)$ (see (3.20)). Note $10<\lambda<\mu<$ 1. In our notation
$f={\rho_{+}^{2}\over l^{2}}\cdot{(x^{2}-1)(x^{2}-\lambda^{2})\over x^{2}},\quad\Delta^{-1}={\rho_{+}^{2}\over l^{2}}c_{1}^{2}\sqrt{1+c_{2}^{2}x^{2}},\quad g=-{\lambda\over lx^{2}}.$ (D.16)D.16( D. 16 )
The solution to the equation (D.7) (for all $x$) is
$\matrix{x^{\star}&=&-{c_{1}^{2}\over 2(1-\lambda^{2})}\pmatrix{-\sqrt{1+c^{2}_{2}}\ln(|1-x^{2}|)+\lambda\sqrt{1+\lambda^{2}c_{2}^{2}}\ln(|\lambda^{2}-x^{2}|)\cr-2|c_{2}|(1-\lambda^{2})\ln\left(1+2c_{2}^{2}x^{2}+2|c_{2}|x\sqrt{1+c_{2}^{2}x^{2}}\right)\cr+\sqrt{1+c_{2}^{2}}\ln\left(1+(1+2c_{2}^{2})x^{2}+2\sqrt{1+c_{2}^{2}}\sqrt{1+c_{2}^{2}x^{2}}x\right)\cr-\lambda\sqrt{1+\lambda^{2}c_{2}^{2}}\ln\left(\lambda^{2}+(1+2\lambda^{2}c_{2}^{2})x^{2}+2\lambda\sqrt{1+\lambda^{2}c_{2}^{2}}\sqrt{1+c_{2}^{2}x^{2}}x\right)},\cr&=&|c_{2}|c_{1}^{2}\ln\left(1+2c_{2}^{2}x^{2}+2|c_{2}|x\sqrt{1+c_{2}^{2}x^{2}}\right)\cr&&+{1\over 2\rho_{+}}\pmatrix{{1\over\kappa_{1}}\ln(|1-x^{2}|)+{1\over\kappa_{2}}\ln(|\lambda^{2}-x^{2}|)\cr-{1\over\kappa_{1}}\ln\left(1+(1+2c_{2}^{2})x^{2}+2\sqrt{1+c_{2}^{2}}\sqrt{1+c_{2}^{2}x^{2}}x\right)\cr-{1\over\kappa_{2}}\ln\left(\lambda^{2}+(1+2\lambda^{2}c_{2}^{2})x^{2}+2\lambda\sqrt{1+\lambda^{2}c_{2}^{2}}\sqrt{1+c_{2}^{2}x^{2}}x\right)},}$ (D.17)D.17( D. 17 )
where $\kappa_{1}$ and $\kappa_{2}$ are given by
$\kappa_{1}={1\over\rho_{+}}\cdot{1-\lambda^{2}\over c_{1}^{2}\sqrt{1+c_{2}^{2}}},\quad\kappa_{2}=-{1\over\rho_{+}}\cdot{1-\lambda^{2}\over\lambda c_{1}^{2}\sqrt{1+\lambda^{2}c_{2}^{2}}},\quad{\kappa_{2}\over\kappa_{1}}<-1.$ (D.18)D.18( D. 18 )
Without loss of generality we can take $c_{2}$ to be positive.
There are two zeros. In the patch around $x=1$ we must take $\kappa=\kappa_{1}$ so that the factor $\Omega^{2}$ is regular at and around $x=1$. We denote this patch by $K_{1}$. It covers the region $x>\lambda$. Note on $K_{1}$ the constant $g_{0}=-\lambda/l$. In the patch around $x=\lambda$ however we must take $\kappa=\kappa_{2}$. This ensures that $\Omega^{2}$ is non-singular at $x=\lambda$. We denote the patch that covers the region $10<x<$ 1 by $K_{2}$. On $K_{2}$ the constant $g_{0}=-1/(l\lambda)$. In the overlapping region $\lambda<x<1$ we identify $K_{1}$ and $K_{2}$. We denote the overlap by $K_{0}$.
In the large $x$ limit we get $x^{\star}=2|c_{2}|c_{1}^{2}\ln x+h(c_{1},c_{2},\lambda)$.28 The BTZ solution is obtained by setting $\mu=1$, i.e., $c_{2}=0$. For $c_{2}=0$, in the large $x$ limit, we instead obatin $x^{\star}=-c_{1}^{2}/x+{\cal O}(x^{-3})$. The function $h$ in particular satisfies $h(c_{1},0,\lambda)=0$. In the large $x$ limit and at constant $\varphi$ the metric (D.3) up to an overall positive factor equals $du^{2}-dv^{2}+[\lambda^{2}/(u^{2}-v^{2})](udv-vdu)^{2}$. Therefore, outgoing and ingoing radial null lines end on future and past null infinities, respectively.
Penrose diagrams for the regions ${\tenrm III}=\{(x,t)|0<x\leq\lambda,t\in{\tenrm I\kern-1.79993ptR}\}$, ${\tenrm II}=\{(x,t)|\lambda<x<1,t\in{\tenrm I\kern-1.79993ptR}\}$ and ${\tenrm I}=\{(x,t)|1\leq x<\infty,t\in{\tenrm I\kern-1.79993ptR}\}$ are given in Fig. 7. The Penrose diagram for the domain III is obtained using $K_{2}$. Similarly, the Penrose diagram for I is obtained using $K_{1}$. The Penrose diagram for II can be obtained using $K_{0}$ or either $K_{1}$ or $K_{2}$. Here the Penrose diagram for II is obtained using $K_{2}$. The Penrose diagram of the whole spacetime is obtained by appropriately gluing together these diagrams. The resulting diagram is given in Fig. 1 in section two.
Fig. 7: The plots depict the Penrose diagrams of the regions I, II and III. Each point represents a circle or the orbit of the isometry $\partial_{\varphi}$. Straight lines at $\pi/4$ radians represent null lines. We used the values $\lambda=0.4,\mu=0.5,A=1$, $\rho_{+}=1$ (in string units), $l=1$ (in string units). For $A=-1$ we simply flips $U\to-U$. Plot a) is the Penrose diagram for III. $x=0$ is mapped to the vertical line $U=\pi/2$. Plot b) is the Penrose diagram for II. It is obtained using $K_{2}$. The lines $u=\pm v$ at $x=\lambda$ are mapped to the lines $U=\pm V$. $x=1$ is mapped to $U\pm V=\pi$. Plot c) is the Penrose diagram for I. Null lines at $x=1$ are mapped to $U=\pm V$. $x=\infty$ is mapped to $U\pm V=\pi$.